### v1

# Dual-scale Enhanced and Cross-generative Consistency Learning for Semi-supervised Polyp Segmentation

Abstract
Automatic polyp segmentation plays a crucial role in the early diagnosis and treatment of colorectal cancer (CRC). However, existing methods heavily rely on fully supervised training, which requires a large amount of labeled data with time-consuming pixel-wise annotations. Moreover, accurately segmenting polyps poses challenges due to variations in shape, size, and location. To address these issues, we propose a novel Dual-scale Enhanced and Cross-generative consistency learning framework for semi-supervised polyp Segmentation (DEC-Seg) from colonoscopy images. First, we propose a Cross-level Feature Aggregation (CFA) module that integrates cross-level adjacent layers to enhance the feature representation ability across different resolutions. To address scale variation, we present a scale-enhanced consistency constraint, which ensures consistency in the segmentation maps generated from the same input image at different scales. This constraint helps handle variations in polyp sizes and improves the robustness of the model. Additionally, we design a scale-aware perturbation consistency scheme to enhance the robustness of the mean teacher model. Furthermore, we propose a cross-generative consistency scheme, in which the original and perturbed images can be reconstructed using cross-segmentation maps. This consistency constraint allows us to mine effective feature representations and boost the segmentation performance. To produce more accurate segmentation maps, we propose a Dual-scale Complementary Fusion (DCF) module that integrates features from two scale-specific decoders operating at different scales. Extensive experimental results on five benchmark datasets demonstrate the effectiveness of our DEC-Seg against other state-of-the-art semi-supervised segmentation approaches. The implementation code will be released at https://github.com/taozh2017/DECSeg.
Index Terms:
Polyp segmentation, semi-supervised learning, scale-enhanced consistency, cross-generative consistency
I Introduction
Colorectal cancer (CRC) is one of the leading causes of death worldwide, often arising from adenomatous polyps. Therefore, early detection and removal of polyps are crucial in reducing the incidence of CRC. Polyp segmentation plays a vital role in identifying and locating polyps in an early stage, enabling clinicians to diagnose and treat them effectively. Currently, various deep learning models [1, 2, 3, 4, 5] have been developed and obtained promising performance in polyp segmentation. However, these methods typically rely on a fully supervised training strategy, which requires a large number of pixel-wise annotations. Unfortunately, annotating medical images often demands highly specialized expertise, making the process time-consuming and costly. As a solution, semi-supervised learning (SSL) provides an effective approach to leverage unlabeled data to enhance the model’s performance.
Recently, various semi-supervised medical image segmentation methods [6, 7, 8, 9, 10, 11, 12, 13, 14] based on deep learning have been developed. Among these methods, the consistent constraint is a widely used strategy, by enhancing that the perturbations on unlabeled data should not significantly vary their outputs or predictions. One of the most representative frameworks is the mean teacher (MT) [15], which designs a perturbation-based consistency loss between the teacher and student models on the unlabeled examples. Inspired by MT, several improved methods focus on designing different perturbation strategies to achieve SSL segmentation. For instance, an uncertainty-aware framework (UA-MT) [10] was proposed to make the student model gradually learn more reliable targets and eliminate unreliable predictions by exploiting the uncertainty information. Chen et al. [7] developed a cross-pseudo supervision strategy, which initializes two identical networks with different weights and encourages high consistency between the predictions of two different networks with different weights for the same input. Luo et al. [16] designed an uncertainty rectified pyramid consistency (URPC) scheme to enable the segmentation model can produce consistent predictions at different scales. Besides, CLCC [17] was proposed to fuse global and patch-wise information by contrastive learning and consistency constraint, which has been applied to polyp segmentation. SLCNet [18] was proposed to utilize shape information for the training by feeding pseudo-labels into the network simultaneously. Despite the achievement made, semi-supervised polyp segmentation remains challenging due to variations in the shape, size, and location of polyps. Therefore, it is crucial to develop effective and robust semi-supervised learning strategies that leverage a large amount of unlabeled data to boost the polyp segmentation performance.
To this end, we propose a novel Dual-scale Enhanced and Cross-generative consistency learning framework for semi-supervised polyp Segmentation (DEC-Seg), which fully exploits multi-scale information from the labeled and unlabeled data to enhance segmentation performance. Specifically, we first propose a Cross-level Feature Aggregation (CFA) module to integrate the adjacent layer features in the encoder, and the cross-level aggregated features are further incorporated into the scale-specific decoder. Then, we present a scale-enhanced consistency strategy to encourage the consistency of segmentation maps from the same inputs with different scales. This consistency can guide our segmentation network to learn more powerful features for handling scale variations. Moreover, a Dual-scale Complementary Fusion (DCF) module is further proposed to filter the effective information from different scale features and fuse them to generate the final prediction map. Besides, we design a scale-aware perturbation consistency to impose on different scale images. Meanwhile, a cross-generative consistency is presented to constrain the original images and perturbed ones can be reconstructed by cross-segmentation maps, which can effectively harness the knowledge from unlabeled data. Finally, we conduct the comparison experiments on five benchmark datasets, and the results demonstrate the superiority of the proposed model over other state-of-the-art semi-supervised segmentation methods.
The main contributions of this paper are listed as follows:
  • We propose a novel semi-supervised polyp segmentation framework, which leverages scale-enhanced consistency and cross-generative consistency to exploit the relationships between labeled and unlabeled data for boosting the segmentation model.
  • A cross-level feature aggregation module is proposed to fuse the cross-level features, which can enhance the representation ability of features within different resolutions.
  • We present a scale-enhanced consistency constraint to reduce the discrepancy between the predicted segmentation maps from different scale inputs. Besides, a cross-generative consistency is designed to constrain that the original and perturbed images can be reconstructed using cross-segmentation maps, which aims to mine effective feature representations and boost the segmentation performance.
  • A dual-scale complementary fusion module is proposed to generate better predictions through the complementation and fusion of different scale information from the two scale-specific decoders.
[…]
Figure 1: Architecture of our DEC-Seg framework. Labeled data are fed to the student model to achieve the segmentation task under supervised learning. Unlabeled data and the corresponding perturbed versions are passed through the student model and teacher model, which are constrained under scale-enhanced consistency, scale-aware perturbation consistency, and cross-generative consistency. See Sec. III for details.
The remainder of this paper is structured as follows. In Section II, we briefly review some related works. In Section III, we give the overview framework and then detail the proposed semi-supervised segmentation model. Then, we present the experimental settings, comparison results, and ablation study in Section IV. Finally, we conclude this paper in Section V.
II Related works
II-A Polyp Segmentation
Recently, numerous studies have proposed different deep learning architectures and techniques to enhance the accuracy of polyp segmentation. The UNet framework, proposed by Ronneberger et al. [19], has significantly advanced the field of medical image segmentation. It incorporates skip connections between the contracting and expanding paths, enabling the network to retain fine-grained spatial information while also capturing high-level contextual features. UNet has demonstrated state-of-the-art performance across various segmentation tasks, including liver and tumor segmentation, cell segmentation, brain MRI segmentation, and polyp segmentation. To further improve the accuracy of UNet-based models, researchers have developed several variants, such as attention UNet [20], UNet++ [1], and ResUNet++ [3]. For example, Zhou et al. [1] presented UNet++ for medical image segmentation which designs a series of nested, dense skip pathways to reduce the semantic gap between the feature maps. Jha et al. [3] developed a ResUNet++ architecture, which takes advantage of residual blocks, squeeze and excitation blocks, atrous spatial pyramidal pooling, and attention blocks. Besides, several polyp segmentation methods [21, 22, 23, 24] have been developed and achieved promising performance.
II-B Semi-Supervised Learning
Existing SSL methods can be categorized into self-training, consistency learning, contrastive learning, and adversarial learning. Self-training [25, 26] methods use the predictions of the fully-supervised algorithm to act as the pseudo labels of the unlabeled data and retrain the model by mixing with the labeled data. Consistency learning methods [16, 27, 15, 7] enforce agreement between different views to improve the model’s performance. By utilizing a consistency loss, these methods ensure that the model’s predictions remain consistent across different representations or augmentations of the same input. Contrastive learning methods [28, 17, 29, 30, 31] obtain better representation learning by contrasting similar features against dissimilar features. Adversarial learning methods [32, 33, 34] employ a minimax game between a generator and a discriminator to train the model. By generating new and realistic samples, these methods enhance the model’s ability to generalize and produce accurate segmentation. Several SSL methods have been developed for the medical image segmentation task. For example, Li et al. [27] proposed a segmentation model with multiple auxiliary decoders and encourage the consistency of the predictions made by the main decoder and the auxiliary decoders. Besides, contrastive learning [28, 17] has been widely used in SSL medical image segmentation. Moreover, adversarial learning [32, 33, 35] has also emerged to improve the robustness of the model. For example, Peiris et al. [32] equipped the network with a critic network to influence the segmentation network to produce the resemble prediction map as ground truth. Lei et al. [33] proposed double discriminators which are used to learn the prior relationship between labeled and unlabeled data. However, these semi-supervised methods do not deeply explore the exploitability and importance of scale information for semi-supervised learning.
II-C Consistency Learning
Consistency learning as one of the most popular methods in semi-supervised learning aims to force the network to learn potential knowledge from different predictions by applying consistency constraints, which may come from different views of the same input from the same network, or may come from the same input from different networks. Currently, several methods [10, 7, 36, 37, 38] based on the mean teacher framework focus on designing different consistent and perturbation strategies to achieve SSL segmentation. For instance, Verma et al. [36] regularized semi-supervised learning by encouraging consistent predictions at interpolations of unlabeled points u1 and u2. Wu et al. [37] proposed three decoders and a different upsampling module in each decoder to amplify the perturbation and impose consistency across the three prediction maps. Zhong et al. [38] proposed a multi-attention tri-branch network (MTNet) that each branch uses a different attention mechanism. Among these methods, the consistent constraint is a widely used strategy, by enhancing that the perturbations on unlabeled data should not significantly vary their outputs.
[…]
Figure 2: The detailed architectures of the encoder (i.e., $E_{s}$ and $E_{t}$ in Fig. 1) and scale-specific decoder (i.e., $D_{s1}$, $D_{s2}$, $D_{t1}$, and $D_{t2}$ in Fig. 1) networks. “Conv” denotes a sequential operation that consists of a $3\times{3}$ convolution, batch normalization, and a ReLU activation, and “Trans-Conv” denotes a transpose convolution.
III Proposed Method
III-A Overview
[…]
III-B Cross-level Feature Aggregation Module
To fully exploit cross-level information and provide more effective supplementary from different convolutional layers, we propose the CFA module to fuse the features from every two adjacent layers. Then, the aggregated feature can be incorporated into our designed decoder, and the feature flow is shown in Fig. 2. Specifically, as shown in Fig. 3, two adjacent features $F_{i}$ and $F_{i+1}$ first proceeded by a $1\times 1$ convolution, and then the two features are concatenated (i.e., $F_{cat}$) to obtain $F_{cat}^{{}^{\prime}}=\mathcal{B}_{conv3\times 3}(F_{cat})$, where $\mathcal{B}_{conv3{\times}3}(\cdot)$ is a sequential operation that consists of a $3\times{3}$ convolution, batch normalization, and a ReLU activation. To learn the cross-level attention-based enhanced feature, we conduct a global average pooling (GAP) on the cascaded feature $F_{cat}^{{}^{\prime}}$ and utilize the point-wise convolution (PWC) [40] to capture channel interactions across different spatial positions. Therefore, we obtain the attention-based weights by
$W=\sigma\big{(}\textup{Conv}_{pwc2}(\zeta(\textup{Conv}_{pwc1}(\mathcal{G}_{ave}(F_{cat}^{{}^{\prime}}))))\big{)},$ (1)
where the kernel sizes of $\textup{Conv}_{pwc1}$ and $\textup{Conv}_{pwc2}$ denote as $\frac{C}{r}\times C\times 1\times 1$ and $C\times\frac{C}{r}\times 1\times 1$ ($C$ is channel size and $r$ is a reduction ratio), respectively. Besides, $\zeta(\cdot)$ and $\sigma(\cdot)$ indicate ReLU and Sigmoid activation functions, respectively, and $\mathcal{G}_{ave}$ is a GAP operation. Next, an element-wise multiplication is utilized to enhance $F_{cat}^{{}^{\prime}}$ with $W$, and then a residual structure is also adopted to fuse the enhanced feature and the original cascaded feature. Finally, we obtain the aggregated feature by
$F_{i}^{\textup{CFA}}=\mathcal{B}_{conv3{\times}3}\big{(}F_{cat}^{{}^{\prime}}\otimes{W}\oplus F_{cat}^{{}^{\prime}}\big{)},$ (2)
where $\oplus$ and $\otimes$ represent element-wise addition and multiplication, respectively. It is worth noting that our CFA module can explore contextual information from the diversity of resolutions to enhance the features’ representation ability.
[…]
Figure 3: Illustrations of the encoder and cross-level feature aggregation module. “GAP” denotes the global average pooling, and “PWC” denotes the point-wise convolution.
III-C Scale-enhanced Consistency
The scale variation is still a great challenge for polyp segmentation. As discussed in [41], reducing the gap between the network outputs from different scale images can help the model learn scale-correlated features. Motivated by this observation, we propose a scale-enhanced consistency scheme to constrain the outputs of different scale images to be closed, in which the different scale features can be refined by each other. Specifically, given unlabeled images $X_{u}$, $X_{u}$ and its downsampled versions $X_{u}^{d}=Down(X_{u})$ are fed to a weight-shared encoder to extract two sets of multi-level features, and then we can obtain the predicted maps $S_{u}$ and $S_{u}^{d}$, respectively. Specifically, the two segmentation maps can be obtained by $S_{u}=D_{s1}(E_{s}(X_{u}))$ and $S_{u}^{d}=D_{s2}(E_{s}(Down(X_{u})))$, where $E_{s}(\cdot)$ is the encoder in the student model, and $Down(\cdot)$ denotes a $1/2$ downsampling operation. $D_{s1}$ and $D_{s2}$ are scale-specific decoders in the student model, which are independent of each other to increase the perturbation. Similarly, for labeled images $X_{l}$, we can obtain the predicted segmentation maps $S_{l}$ and $S_{l}^{d}$ from $X_{l}$ and its downsampled versions $X_{l}^{d}$ by $S_{l}=D_{s1}(E_{s}(X_{l}))$ and $S_{l}^{d}=D_{s2}(E_{s}(Down(X_{l})))$, respectively. Therefore, we form a scale-enhanced consistency loss ($\mathcal{L}_{SC}$) to enforce $S$ and $S^{d}$ to be consistent, which is defined by
$\mathcal{L}_{SC}=\mathcal{L}_{mse}(S_{u},S_{u}^{d})+\mathcal{L}_{mse}(S_{l},S_{l}^{d}),$ (3)
where $\mathcal{L}_{mse}$ denotes the widely used mean square error loss. Finally, the supervised loss is formulated by
$\begin{split}\mathcal{L}_{S}=&\mathcal{L}_{sup}(S_{l},Y_{l})+\mathcal{L}_{sup}(S_{l}^{d},Down(Y_{l})).\end{split}$ (4)
III-D Dual-scale Complementary Fusion Module
[…]
$h_{i}^{dcf}=h_{rec}^{1}\otimes\alpha_{1}+h_{rec}^{2}\otimes\alpha_{2}.$ (5)
Subsequently, the fused features $h_{i}^{dcf}$ will be passed into the fused decoder. Specifically, $h_{5}^{dcf}$ is first passed through a transpose convolutional layer to have the same resolution with $h_{4}^{dcf}$. Then, the two features are concatenated and fed into two convolutional layers. Repeating the above process, we will obtain the final segmentation maps $S_{l}^{f}$ for labeled data. Therefore, the supervised loss in Eq. (4) can be reformulated as follows:
$\begin{split}\mathcal{L}_{S}=&\mathcal{L}_{sup}(S_{l},Y_{l})+\\ &\mathcal{L}_{sup}(S_{l}^{d},Down(Y_{l}))+\mathcal{L}_{sup}(S_{l}^{f},Y_{l}).\end{split}$ (6)
[…]
Figure 4: Illustration of (a) the architecture of the fused decoder (i.e., ${D}_{s}^{f}$ and ${D}_{t}^{f}$ in Fig. 1) and (b) the proposed dual-scale complementary fusion module.
[…]
Figure 5: The detailed architecture of the generative network (i.e., $G_{1}$ and $G_{2}$). “Conv” denotes a sequential operation that consists of a $3\times{3}$ convolution, batch normalization, and a ReLU activation, and “Trans-Conv” denotes a transpose convolution.
III-E Scale-aware Perturbation and Cross-generative Consistency
Following previous works [35, 33, 27, 42, 43], to improve the robustness of the MT model, $X_{u}$ and its perturbed version $X_{u}^{p}$ are passed through both student and teacher models, respectively, to encourage the outputs to be closed. In this study, we design a scale-aware perturbation consistency. Specifically, we impose the perturbation consistency on different scale images, i.e., $S_{u}^{p}=D_{t1}(E_{t}(\mathcal{P}(X_{u})))$ and $S_{u}^{dp}=D_{t2}(E_{t}(\mathcal{P}(Down(X_{u}))))$, where a random gamma operation is adopted as the perturbation operation (i.e., $\mathcal{P}$). Meanwhile, we also obtain ${S}_{u}^{fp}$ that its constraint is implemented by the MT framework, due to the particularity of ${S}_{u}^{fp}$. Thus, the scale-aware perturbation consistency scheme can be optimized by minimizing $\mathcal{L}_{SPC}$, which is defined by
$\mathcal{L}_{SPC}=\mathcal{L}_{mse}(S_{u},S_{u}^{p})+\mathcal{L}_{mse}(S_{u}^{d},S_{u}^{dp})+\mathcal{L}_{mse}(S_{u}^{f},S_{u}^{fp}).$ (7)
Further, to utilize the rich information contained in different feature maps, we propose a perturbation-based cross-generative consistency constraint to enhance the feature mining ability of our segmentation network. Therefore, we design two generative networks $G_{1}$ and $G_{2}$ (The detailed structure is shown in Fig. 5) for image reconstruction. The segmentation maps $S_{u}^{f}$ and $S_{u}^{fp}$ obtained from the aggregation of the two scale information are fed into $G_{2}$ and $G_{1}$, respectively, and then we can obtain two reconstructed images $\hat{X}_{u}=G_{1}(S_{u}^{fp})$ and $\hat{X}_{u}^{p}=G_{2}(S_{u}^{f})$. Finally, we form a cross-generative consistency loss ($\mathcal{L}_{CC}$), which is given by
$\mathcal{L}_{CC}=\mathcal{L}_{mse}(\hat{X}_{u},X_{u})+\mathcal{L}_{mse}(\hat{X}_{u}^{p},X_{u}^{p}).$ (8)
It is worth noting that $\mathcal{L}_{CC}$ can enforce $\hat{X}_{u}$ and $X_{u}$ to be consistent and $\hat{X}_{u}^{p}$ and $X_{u}^{p}$ to be closed. As a result, such cross-generation can further improve the robustness of our MT-based model to learn powerful features and fully harness the knowledge from unlabeled data to boost the proposed segmentation model.
TABLE I: Quantitative results on CVC-ClinicDB and Kvasir datasets with $10\%$ or $30\%$ labeled data.
Methods 10% labeled 30% labeled
mDice $\uparrow$ mIoU $\uparrow$ $F_{\beta}^{w}$uparrow $S_{\alpha}$uparrow MAE $\downarrow$ mDice $\uparrow$ mIoU $\uparrow$ $F_{\beta}^{w}$uparrow $S_{\alpha}$uparrow MAE $\downarrow$
CVC-ClinicDB Baseline 0.767 0.689 0.756 0.847 0.024 0.812 0.751 0.806 0.873 0.021
MT [15] 0.785 0.711 0.763 0.855 0.025 0.839 0.781 0.829 0.892 0.020
DAN [34] 0.789 0.718 0.779 0.862 0.024 0.838 0.781 0.835 0.890 0.018
UA-MT [10] 0.798 0.729 0.782 0.869 0.022 0.831 0.762 0.817 0.879 0.026
URPC [16] 0.794 0.723 0.783 0.864 0.022 0.829 0.767 0.825 0.880 0.022
Duo-SegNet [32] 0.765 0.685 0.745 0.840 0.037 0.755 0.678 0.739 0.832 0.038
CLCC [17] 0.794 0.720 0.786 0.859 0.027 0.840 0.786 0.840 0.891 0.019
SLC-Net [18] 0.752 0.689 0.735 0.844 0.027 0.835 0.774 0.830 0.888 0.020
MC-Net+ [37] 0.767 0.702 0.751 0.856 0.025 0.845 0.783 0.844 0.891 0.020
CDMA [38] 0.759 0.676 0.750 0.837 0.025 0.803 0.728 0.802 0.863 0.024
BCP [44] 0.835 0.769 0.828 0.887 0.018 0.846 0.782 0.847 0.894 0.020
DEC-Seg (ours) 0.830 0.772 0.836 0.888 0.014 0.852 0.798 0.853 0.897 0.019
Kvasir Baseline 0.786 0.701 0.756 0.831 0.059 0.843 0.782 0.829 0.877 0.037
MT [15] 0.833 0.755 0.818 0.858 0.051 0.868 0.799 0.853 0.881 0.037
DAN [34] 0.836 0.758 0.820 0.858 0.049 0.867 0.803 0.851 0.886 0.035
UA-MT [10] 0.824 0.753 0.808 0.859 0.051 0.858 0.787 0.838 0.881 0.039
URPC [16] 0.821 0.740 0.803 0.853 0.050 0.854 0.789 0.843 0.879 0.037
Duo-SegNet [32] 0.795 0.711 0.785 0.838 0.061 0.851 0.772 0.843 0.878 0.043
CLCC [17] 0.806 0.724 0.796 0.844 0.052 0.864 0.804 0.853 0.892 0.034
SLC-Net [18] 0.840 0.773 0.830 0.868 0.042 0.867 0.805 0.856 0.890 0.034
MC-Net+ [37] 0.817 0.735 0.807 848 0.051 0.831 0.765 0.816 0.862 0.044
CDMA [38] 0.786 0.695 0.769 0.826 0.056 0.839 0.758 0.826 0.862 0.044
BCP [44] 0.854 0.789 0.850 0.876 0.042 0.886 0.820 0.876 0.898 0.029
DEC-Seg (ours) 0.862 0.794 0.853 0.880 0.040 0.882 0.822 0.871 0.899 0.033
III-F Loss Function and Optimization
Inspired by [45], the supervised loss for labeled data is formulated by $\mathcal{L}_{sup}=\mathcal{L}_{IoU}^{w}+\mathcal{L}_{BCE}^{w}$, where $\mathcal{L}_{IoU}^{w}$ and $\mathcal{L}_{BCE}^{w}$ denote the weighted IoU loss and binary cross-entropy (BCE) loss, respectively. Finally, the total loss can be given as follows:
$\mathcal{L}_{total}=\mathcal{L}_{S}+\mathcal{L}_{SPC}+\alpha\mathcal{L}_{SC}+\beta\mathcal{L}_{CC},$ (9)
where the hyper-parameter $\alpha$ is set to $1.5$ and $\beta$ is set to $3$. In addition, the parameters $\theta^{{}^{\prime}}$ of the teacher network can be updated based on the parameters $\theta$ of the student network using an exponential moving average (EMA) strategy. Thus, at the $t$-th training step, the parameters $\theta^{{}^{\prime}}$ can be updated by $\theta^{{}^{\prime}}_{t}=\gamma\theta^{{}^{\prime}}_{t-1}+(1-\gamma)\theta_{t}$, where $\gamma$ is used to control the EMA decay, and it is set to $0.99$.
IV EXPERIMENTS AND RESULTS
IV-A Experimental Setup
IV-A1 Datasets and Training Settings
We carry out the comparison experiments on five public datasets, i.e., ETIS [46], CVC-ClinicDB [47], EndoScene-CVC300 [48], CVC-ColonDB [49], and Kvasir [50]. Following the setting in [2], $1450$ images are randomly selected from the two datasets (Kvasir and CVC-ClinicDB) for the training set, and the remaining images from the two datasets and the other three datasets (EndoScene-CVC300, ETIS, and CVC-ColonDB) to form the testing set. Besides, $10\%$ ($145$ images) or $30\%$ ($435$ images) from the training set are used as the labeled data, while the remaining images are adopted as the unlabeled data.
IV-A2 Implementation Details and Evaluation Metrics
The proposed segmentation framework is implemented in PyTorch and trained on one NVIDIA GeForce RTX3090 GPU with an Adam optimizer. All inputs are uniformly rescaled to $352\times 352$. Our model converges over $10000$ iterations with a batch size of $8$ including 4 labeled samples and 4 unlabeled samples, and the learning rate is set to $1e$-$4$.
To evaluate the effectiveness, we employ five commonly adopted metrics [2, 3], namely mean Dice (mDice), mean IoU (mIoU), $F_{\beta}^{w}$, $S_{\alpha}$, and mean absolute error (MAE).
TABLE II: Quantitative results on three unseen datasets (EndoScene-CVC300, CVC-ColonDB, and ETIS) with $10\%$ or $30\%$ labeled data.
Methods 10% labeled 30% labeled
mDice $\uparrow$ mIoU $\uparrow$ $F_{\beta}^{w}$uparrow $S_{\alpha}$uparrow MAE $\downarrow$ mDice $\uparrow$ mIoU $\uparrow$ $F_{\beta}^{w}$uparrow $S_{\alpha}$uparrow MAE $\downarrow$
EndoScene-CVC300 Baseline 0.800 0.708 0.784 0.861 0.013 0.827 0.760 0.812 0.895 0.008
MT [15] 0.858 0.779 0.837 0.910 0.010 0.873 0.796 0.845 0.916 0.009
DAN [34] 0.820 0.733 0.797 0.885 0.014 0.874 0.800 0.847 0.917 0.008
UA-MT [10] 0.859 0.780 0.829 0.910 0.010 0.837 0.754 0.795 0.895 0.012
URPC [16] 0.841 0.753 0.812 0.896 0.012 0.872 0.800 0.859 0.920 0.007
Duo-SegNet [32] 0.841 0.761 0.811 0.894 0.011 0.843 0.758 0.803 0.902 0.012
CLCC [17] 0.748 0.674 0.724 0.845 0.015 0.826 0.755 0.811 0.891 0.008
SLC-Net [18] 0.856 0.777 0.827 0.907 0.012 0.869 0.796 0.849 0.915 0.010
MC-Net+ [37] 0.783 0.711 0.755 0.872 0.014 0.845 0.783 0.844 0.891 0.020
CDMA [38] 0.791 0.685 0.765 0.846 0.013 0.803 0.728 0.802 0.863 0.024
BCP [44] 0.856 0.770 0.835 0.901 0.010 0.846 0.782 0.847 0.894 0.020
DEC-Seg (ours) 0.860 0.781 0.831 0.914 0.011 0.895 0.824 0.875 0.927 0.006
CVC-ColonDB Baseline 0.531 0.453 0.526 0.712 0.052 0.551 0.489 0.543 0.732 0.049
MT [15] 0.610 0.531 0.597 0.759 0.044 0.668 0.587 0.652 0.788 0.041
DAN [34] 0.599 0.515 0.593 0.748 0.049 0.629 0.551 0.613 0.766 0.043
UA-MT [10] 0.583 0.507 0.576 0.745 0.050 0.666 0.587 0.641 0.785 0.048
URPC [16] 0.587 0.507 0.574 0.745 0.048 0.621 0.551 0.612 0.768 0.045
Duo-SegNet [32] 0.644 0.550 0.613 0.763 0.058 0.658 0.569 0.627 0.773 0.057
CLCC [17] 0.538 0.473 0.533 0.720 0.053 0.564 0.505 0.562 0.738 0.050
SLC-Net [18] 0.595 0.525 0.582 0.755 0.044 0.653 0.582 0.644 0.783 0.041
MC-Net+ [37] 0.562 0.486 0.543 0.725 0.055 0.589 0.518 0.578 0.745 0.050
CDMA [38] 0.507 0.419 0.494 0.691 0.054 0.624 0.536 0.609 0.756 0.047
BCP [44] 0.642 0.563 0.635 0.773 0.046 0.656 0.580 0.651 0.778 0.046
DEC-Seg (ours) 0.650 0.575 0.639 0.782 0.042 0.691 0.616 0.678 0.798 0.042
ETIS Baseline 0.458 0.392 0.433 0.693 0.024 0.486 0.432 0.469 0.715 0.021
MT [15] 0.468 0.395 0.434 0.694 0.036 0.595 0.519 0.552 0.761 0.035
DAN [34] 0.473 0.395 0.440 0.700 0.031 0.528 0.452 0.485 0.729 0.045
UA-MT [10] 0.469 0.399 0.444 0.695 0.027 0.603 0.515 0.547 0.761 0.032
URPC [16] 0.459 0.391 0.436 0.691 0.028 0.601 0.523 0.572 0.765 0.023
Duo-SegNet [32] 0.450 0.382 0.407 0.669 0.061 0.578 0.500 0.531 0.750 0.049
CLCC [17] 0.409 0.346 0.395 0.668 0.031 0.474 0.422 0.465 0.707 0.020
SLC-Net [18] 0.431 0.374 0.409 0.682 0.030 0.566 0.503 0.543 0.753 0.024
MC-Net+ [37] 0.439 0.369 0.403 0.684 0.035 0.535 0.470 0.505 0.735 0.026
CDMA [38] 0.309 0.254 0.296 0.611 0.031 0.531 0.446 0.494 0.724 0.024
BCP [44] 0.571 0.502 0.550 0.756 0.023 0.557 0.497 0.541 0.754 0.020
DEC-Seg (ours) 0.615 0.540 0.582 0.775 0.019 0.683 0.609 0.658 0.814 0.014
IV-B Comparison with State-of-the-arts
We compare the proposed DEC-Seg with ten state-of-the-art semi-supervised segmentation methods, i.e., MT [15], DAN [34], UA-MT [10], URPC [16], Duo-SegNet [32], CLCC [17], SLC-Net [18], MC-Net+ [37], CDMA [38] and BCP [44]. For a fair comparison, we change the backbones of all compared methods to “Res2Net”, and then train all compared methods without any data augmentation. Moreover, we adopt a basic U-Net [19] model with Res2Net as the backbone of the encoder, denoted as “Baseline”.
IV-B1 Quantitative Results
Quantitative comparisons are reported in Tab. I and Tab. II. To validate the learning ability of the two seen training datasets (CVC-ClinicDB and Kvasir), we can be seen from Tab. I that while most methods learn well, our method performs better than other compared methods, both on $10\%$ and $30\%$ of labeled data. In addition, due to the additional unlabeled data, our method improves mDice and mIoU by $8.2\%$ and $12.0\%$ compared to baseline using only $10\%$ labeled data, and by $4.9\%$ and $6.3\%$ compared to baseline using only $30\%$ labeled data on CVC-ClinicDB, which indicate that our DEC-Seg can fully leverage unlabeled data to improve the polyp segmentation performance.
Additionally, we report the comparison results on three unseen datasets (CVC-300, ETIS, and CVC-ColonDB) from Tab. II to verify the generalization ability of our DEC-Seg. Because the polyp in the CVC-300 dataset is similar to the training set and relatively simple, the metrics are also similar to CVC-ClinicDB and Kvisir datasets. Compared with SLC-Net, which performs relatively stable on this dataset, our method achieves $3.0\%$ and $3.5\%$ improvements in terms of mDice and mIoU with $30\%$ labeled data. The CVC-ColonDB and ETIS datasets are more difficult, as polyps are with large variations in size. As reported in Tab. II, comparing the BCP method on the CVC-ColonDB dataset using $10\%$ labeled data, our method achieves $1.2\%$ and $2.1\%$ improvements on mDice and mIoU, respectively. Besides, our method achieves $5.3\%$ and $6.2\%$ improvements over the SOTA semi-supervised segmentation method (BCP) in terms of mDice and mIoU with the $30\%$ labeled data. Similarly, on the ETIS dataset, our performance has a more significant superiority over other methods. The main reason is that our exploration and utilization of scale information make the model more robust to scale variation, which is able to identify small polyps that cannot be identified by other methods, and can also refine the edges of large polyps. Overall, the results indicate that our DEC-Seg has a better generalization ability and effectively utilizes unlabeled data to boost polyp segmentation.
[…]
Figure 6: Qualitative results of our model and other ten compared semi-supervised segmentation methods.
IV-B2 Qualitative Results
Fig. 6 shows some of the visualization results of the test examples under $30\%$ labeled data. It can be seen that our method can accurately locate and segment polyps under different challenging factors. For example, in the first two rows of Fig. 6, the polyps have very small sizes. It can be clearly seen that Duo-SegNet recognizes the polyp incorrectly, while UAMT, CLCC, MC-Net+, and CDMA methods even fail to locate polyps. Other methods only locate some fragments of polyps, while our method accurately and completely segments polyps. In the $3^{rd}$ and $4^{th}$ rows, the polyps are visually embedded in their surrounding mucosa, thus it is very difficult to accurately locate and segment these polyps. From the results, our method performs better than other comparison methods to accurately segment polyps. In the $5^{th}$ and $6^{th}$ rows, the polyps have relatively large sizes, making it is challenging to complete locate the polyps. In this case, some methods (e.g., URPC, Duo-SegNet, MC-Net+, and BCP) produce several regions with over-segmented fragments, while our method can obtain promising segmentation results and produce fine details of the boundary. This is mainly because our method makes use of multi-scale information and cross-generative consistency to learn more powerful feature representations, which are helpful to boost the segmentation performance. In addition, we also show some segmentation results containing multiple polyps (see the $7^{th}$ and $8^{th}$ rows). Compared with other methods, the segmentation of multiple polyps by our method is more complete and accurate. It can be also observed that our method can effectively locate and segment polyps under different challenging factors, such as scale variation, homogeneous regions, non-sharp boundaries, and multiple polyps.
TABLE III: Ablative results on the Kvasir and CVC-ColonDB datasets with $30\%$ labeled data.
Settings Kvasir CVC-ColonDB
mDice mIoU $F_{\beta}^{w}$ $S_{\alpha}$ mDice mIoU $F_{\beta}^{w}$ $S_{\alpha}$
(No.1) Baseline 0.843 0.782 0.829 0.877 0.551 0.489 0.543 0.732
(No.2) Baseline + SC 0.859 0.801 0.850 0.886 0.600 0.538 0.593 0.757
(No.3) Baseline + SC + SPC 0.865 0.810 0.856 0.891 0.611 0.548 0.602 0.765
(No.4) Baseline + SC + SPC + CC 0.875 0.815 0.864 0.892 0.635 0.565 0.627 0.773
(No.5) Baseline + SC + SPC + CC + CFA 0.876 0.815 0.867 0.892 0.654 0.580 0.649 0.778
(No.6) Baseline + SC + SPC + CC + CFA + DCF (ours) 0.882 0.822 0.871 0.899 0.691 0.616 0.678 0.798
IV-C Ablation Study
To verify the effectiveness of each key component in DEC-Seg, we conduct ablation studies with $30\%$ labeled data on Kvasir and CVC-ColonDB datasets. The ablative results are shown in Tab. III, where “Baseline” denotes the base segmentation framework only using the labeled data.
Effectiveness of SC. To investigate the importance of scale-enhanced consistency (SC), we add the DC loss to encourage the consistency of predictions from the same inputs with different scales. From Tab. III, we observe that No.2 (Baseline + SC) outperforms No.1 and obtains an $8.9\%$ improvement in mean Dice. This result indicates that scale-enhanced consistency is very helpful to improve polyp segmentation performance and is robust to scale variation.
Effectiveness of SPC and CC. We further study the contributions of scale-aware perturbation consistency (SPC) and cross-generative consistency (CC). As shown in Tab. III, No.3 performs better than No.2, indicating the effectiveness of scale-aware perturbation consistency. Besides, it can be seen that No.4 improves the No.3 performance on the Kvasir dataset, as the mean Dice is improved from $0.865$ to $0.875$. Therefore, these improvements indicate that introducing cross-generative consistency loss can help accurately segment polyp tissues in the learning of details and textures.
Effectiveness of CFA: We then investigate the importance of the proposed CFA module. To achieve this, we add the CFA on the basis of No.4, by aggregating the features of adjacent layers in the encoder before feeding into the decoder. As reported in Tab. III, it can be clearly seen that No.5 has a significant improvement over No.4 on the CVC-ColonDB dataset with a large scale variation, which proves that CFA can effectively capture scale information and improve segmentation performance.
Effectiveness of DCF. As shown in Tab. III, No.6 (using the proposed DCF module) outperforms No.5 on two datasets. This indicates that fusion of multi-scale features can further improve the segmentation performance. In addition, to further validate the effectiveness of the dual-scale complement fusion strategy in the proposed DCF module, we construct a “Basic” strategy, which conducts a concatenation operation followed by two convolution layers to integrate the two features from different scales. The comparison results are shown in Tab. IV. From Tab. IV, it can be observed that our DCF performs better than the “Basic” strategy, indicating the effectiveness of the designed dual-scale complementary fusion module. Moreover, we visualize the segmentation results by using three different decoders, i.e., the scale-specific and fused decoders, and the comparison results are shown in Fig. 7. It can be observed that our method can not accurately locate the boundaries of the polyps when only using the original scale features or the downsampled scale features. However, we integrate the features from the two scales and then propagate them into the fused decoder, which can produce more accurate segmentation maps (as shown in Fig. 7 (c)). This further validates that the fusion of the features from different scales is helpful for improving the segmentation performance.
TABLE IV: Ablative results of our DCF module and a basic fusion strategy on the Kvasir and CVC-ColonDB datasets. The labeled ratio is set to $30\%$.
Settings Kvasir CVC-ColonDB
mDice mIoU $F_{\beta}^{w}$ mDice mIoU $F_{\beta}^{w}$
Basic 0.878 0.810 0.866 0.684 0.604 0.676
DCF 0.882 0.822 0.871 0.691 0.616 0.678
[…]
Figure 7: Visualization results of the three decoders. (a) predictions of the decoder (i.e., $D_{s1}$, $D_{t1}$) using the original scale features; (b) predictions of the decoder (i.e., $D_{s2}$, $D_{t2}$) using the downsampled scale features; (c) predictions of the fused decoders (i.e., $D_{s}^{f}$, $D_{t}^{f}$) using the multi-scale integrated features. The red and blue lines denote the ground truth and predictions, respectively.
V Conclusion
In this paper, we have presented a novel semi-supervised learning framework (DEC-Seg) for polyp segmentation with colonoscopy images. The proposed cross-level feature aggregation module integrates the adjacent features from different resolutions, to enhance the features’ representation ability. Then, a scale-enhanced consistency is proposed to handle scale variation and learn more scale-aware features. Meanwhile, we design the fused decoders and a dual-scale complementary fusion module to aggregate the features from the scale-specific decoders. Moreover, multiple consistency strategies, i.e., scale-aware perturbation consistency and cross-generative consistency, are presented to enhance the learning process and fully leverage unlabeled data to boost the segmentation performance. Further, we propose the dual-scale complementary fusion module to integrate the features from two scale-specific decoders, which are passed through the fused decoder to produce the final segmentation maps. Experimental results on the five benchmark datasets show that our DEC-Seg is superior to state-of-the-art semi-supervised segmentation methods.

### v2

@@
-Dual-scale Enhanced and Cross-generative Consistency Learning for Semi-supervised Polyp Segmentation
+Dual-scale Enhanced and Cross-generative Consistency Learning for Semi-supervised Medical Image Segmentation
 Abstract
-Automatic polyp segmentation plays a crucial role in the early diagnosis and treatment of colorectal cancer (CRC).
+Medical image segmentation plays a crucial role in computer-aided diagnosis.
 However, existing methods heavily rely on fully supervised training, which requires a large amount of labeled data with time-consuming pixel-wise annotations.
-Moreover, accurately segmenting polyps poses challenges due to variations in shape, size, and location.
-To address these issues, we propose a novel Dual-scale Enhanced and Cross-generative consistency learning framework for semi-supervised polyp Segmentation (DEC-Seg) from colonoscopy images.
+Moreover, accurately segmenting lesions poses challenges due to variations in shape, size, and location.
+To address these issues, we propose a novel Dual-scale Enhanced and Cross-generative consistency learning framework for semi-supervised medical image Segmentation (DEC-Seg).
 First, we propose a Cross-level Feature Aggregation (CFA) module that integrates cross-level adjacent layers to enhance the feature representation ability across different resolutions.
 To address scale variation, we present a scale-enhanced consistency constraint, which ensures consistency in the segmentation maps generated from the same input image at different scales.
-This constraint helps handle variations in polyp sizes and improves the robustness of the model.
-Additionally, we design a scale-aware perturbation consistency scheme to enhance the robustness of the mean teacher model.
+This constraint helps handle variations in lesion sizes and improves the robustness of the model.
 Furthermore, we propose a cross-generative consistency scheme, in which the original and perturbed images can be reconstructed using cross-segmentation maps.
 This consistency constraint allows us to mine effective feature representations and boost the segmentation performance.
-To produce more accurate segmentation maps, we propose a Dual-scale Complementary Fusion (DCF) module that integrates features from two scale-specific decoders operating at different scales.
-Extensive experimental results on five benchmark datasets demonstrate the effectiveness of our DEC-Seg against other state-of-the-art semi-supervised segmentation approaches.
+To further exploit the scale information, we propose a Dual-scale Complementary Fusion (DCF) module that integrates features from two scale-specific decoders operating at different scales to help produce more accurate segmentation maps.
+Extensive experimental results on multiple medical segmentation tasks (polyp, skin lesion, and brain glioma) demonstrate the effectiveness of our DEC-Seg against other state-of-the-art semi-supervised segmentation approaches.
 The implementation code will be released at https://github.com/taozh2017/DECSeg.
 Index Terms:
-Polyp segmentation, semi-supervised learning, scale-enhanced consistency, cross-generative consistency
+Medical image segmentation, semi-supervised learning, scale-enhanced consistency, cross-generative consistency.
 I Introduction
-Colorectal cancer (CRC) is one of the leading causes of death worldwide, often arising from adenomatous polyps.
-Therefore, early detection and removal of polyps are crucial in reducing the incidence of CRC.
-Polyp segmentation plays a vital role in identifying and locating polyps in an early stage, enabling clinicians to diagnose and treat them effectively.
-Currently, various deep learning models [1, 2, 3, 4, 5] have been developed and obtained promising performance in polyp segmentation.
-However, these methods typically rely on a fully supervised training strategy, which requires a large number of pixel-wise annotations.
-Unfortunately, annotating medical images often demands highly specialized expertise, making the process time-consuming and costly.
-As a solution, semi-supervised learning (SSL) provides an effective approach to leverage unlabeled data to enhance the model’s performance.
-Recently, various semi-supervised medical image segmentation methods [6, 7, 8, 9, 10, 11, 12, 13, 14] based on deep learning have been developed.
-Among these methods, the consistent constraint is a widely used strategy, by enhancing that the perturbations on unlabeled data should not significantly vary their outputs or predictions.
+Accurate segmentation of organs or lesions from medical images is a crucial task within the realm of medical image processing and analysis.
+The successful accomplishment of this task holds immense significance for disease diagnosis, organ localization, and lesion segmentation.
+In recent years, several deep learning models [1, 2, 3, 4] have been developed and have shown promising performance in medical image segmentation.
+However, these models often rely on a fully supervised training strategy, which necessitates a substantial number of pixel-wise annotations.
+Unfortunately, the annotation process for medical images is both time-consuming and costly due to the need for specialized expertise.
+To address these challenges, researchers have turned to semi-supervised learning (SSL) as an effective approach to leverage unlabeled data and enhance the performance of segmentation models [5, 6].
+By utilizing SSL techniques, models can exploit the abundance of available unlabeled data in medical imaging datasets, reducing their dependence on fully annotated data.
+This approach not only alleviates the burden of extensive annotation requirements but also has the potential to improve the model’s accuracy in medical image segmentation tasks.
+Recently, various semi-supervised medical image segmentation methods [7, 8, 9] based on deep learning have been developed.
+Among these methods, the consistent constraint is a widely used strategy, by ensuring that the perturbations on unlabeled data should not significantly vary their outputs or predictions.
 One of the most representative frameworks is the mean teacher (MT) [10], which designs a perturbation-based consistency loss between the teacher and student models on the unlabeled examples.
@@
 For instance, an uncertainty-aware framework (UA-MT) [11] was proposed to make the student model gradually learn more reliable targets and eliminate unreliable predictions by exploiting the uncertainty information.
+Verma et al.
+[12] proposed an interpolation consistency training framework, which constrains the prediction at an interpolation of unlabeled points to be consistent with that of the predictions at those points.
 Chen et al.
@@
 [14] designed an uncertainty rectified pyramid consistency (URPC) scheme to enable the segmentation model can produce consistent predictions at different scales.
-Besides, CLCC [17] was proposed to fuse global and patch-wise information by contrastive learning and consistency constraint, which has been applied to polyp segmentation.
+Besides, CLCC [15] was proposed to fuse global and patch-wise information by contrastive learning and consistency constraint.
 SLCNet [16] was proposed to utilize shape information for the training by feeding pseudo-labels into the network simultaneously.
-Despite the achievement made, semi-supervised polyp segmentation remains challenging due to variations in the shape, size, and location of polyps.
-Therefore, it is crucial to develop effective and robust semi-supervised learning strategies that leverage a large amount of unlabeled data to boost the polyp segmentation performance.
-To this end, we propose a novel Dual-scale Enhanced and Cross-generative consistency learning framework for semi-supervised polyp Segmentation (DEC-Seg), which fully exploits multi-scale information from the labeled and unlabeled data to enhance segmentation performance.
+Despite the achievement made, semi-supervised medical image segmentation remains challenging due to variations in the shape, size, and location of lesions.
+Therefore, it is crucial to develop effective and robust semi-supervised learning strategies that leverage a large amount of unlabeled data to boost the medical image segmentation performance.
+To this end, we propose a novel Dual-scale Enhanced and Cross-generative consistency learning framework for semi-supervised medical image Segmentation (DEC-Seg), which fully exploits multi-scale information from the labeled and unlabeled data to enhance segmentation performance.
 Specifically, we first propose a Cross-level Feature Aggregation (CFA) module to integrate the adjacent layer features in the encoder, and the cross-level aggregated features are further incorporated into the scale-specific decoder.
@@
 This consistency can guide our segmentation network to learn more powerful features for handling scale variations.
-Moreover, a Dual-scale Complementary Fusion (DCF) module is further proposed to filter the effective information from different scale features and fuse them to generate the final prediction map.
-Besides, we design a scale-aware perturbation consistency to impose on different scale images.
+Moreover, a Dual-scale Complementary Fusion (DCF) module is further proposed to filter the effective information from different scale features and fuse them to boost segmentation performance.
 Meanwhile, a cross-generative consistency is presented to constrain the original images and perturbed ones can be reconstructed by cross-segmentation maps, which can effectively harness the knowledge from unlabeled data.
-Finally, we conduct the comparison experiments on five benchmark datasets, and the results demonstrate the superiority of the proposed model over other state-of-the-art semi-supervised segmentation methods.
+Finally, we conduct the comparison experiments on three medical image segmentation tasks, and the results demonstrate the superiority of the proposed model over other state-of-the-art semi-supervised segmentation methods.
 The main contributions of this paper are listed as follows:
-  • We propose a novel semi-supervised polyp segmentation framework, which leverages scale-enhanced consistency and cross-generative consistency to exploit the relationships between labeled and unlabeled data for boosting the segmentation model.
+  • We propose a novel semi-supervised medical image segmentation framework, which leverages scale-enhanced consistency and cross-generative consistency to exploit the relationships between labeled and unlabeled data for boosting the segmentation model.
   • A cross-level feature aggregation module is proposed to fuse the cross-level features, which can enhance the representation ability of features within different resolutions.
@@
 Besides, a cross-generative consistency is designed to constrain that the original and perturbed images can be reconstructed using cross-segmentation maps, which aims to mine effective feature representations and boost the segmentation performance.
-  • A dual-scale complementary fusion module is proposed to generate better predictions through the complementation and fusion of different scale information from the two scale-specific decoders.
+  • A dual-scale complementary fusion module is proposed to benifit the better prediction generation through the complementation and fusion of different scale information from the two scale-specific decoders.
 […]
 Figure 1: Architecture of our DEC-Seg framework.
-Labeled data are fed to the student model to achieve the segmentation task under supervised learning.
-Unlabeled data and the corresponding perturbed versions are passed through the student model and teacher model, which are constrained under scale-enhanced consistency, scale-aware perturbation consistency, and cross-generative consistency.
+Labeled data are fed to the model to achieve the segmentation task under supervised learning.
+Unlabeled data and the corresponding perturbed versions are passed through the model, which is constrained under scale-enhanced consistency, scale-aware perturbation consistency, and cross-generative consistency.
 See Sec.
@@
 The remainder of this paper is structured as follows.
-In Section II, we briefly review some related works.
-In Section III, we give the overview framework and then detail the proposed semi-supervised segmentation model.
-Then, we present the experimental settings, comparison results, and ablation study in Section IV.
+In Section II, we briefly review some related works to our model, consisting of medical image segmentation, semi-supervised learning, and consistency learning.
+In Section III, we give the overview framework and then provide the details of the proposed semi-supervised segmentation model.
+Then, we present the used datasets, experimental settings, comparison results, and ablation study in Section IV.
 Finally, we conclude this paper in Section V.
 II Related Works
-II-A Polyp Segmentation
-Recently, numerous studies have proposed different deep learning architectures and techniques to enhance the accuracy of polyp segmentation.
+II-A Medical Image Segmentation
+Recently, numerous studies have proposed different deep learning architectures and techniques to enhance the accuracy of medical image segmentation.
 The UNet framework, proposed by Ronneberger et al.
@@
 It incorporates skip connections between the contracting and expanding paths, enabling the network to retain fine-grained spatial information while also capturing high-level contextual features.
-UNet has demonstrated state-of-the-art performance across various segmentation tasks, including liver and tumor segmentation, cell segmentation, brain MRI segmentation, and polyp segmentation.
+UNet has demonstrated state-of-the-art performance across various segmentation tasks, including liver and tumor segmentation, cell segmentation, brain tumor segmentation, and polyp segmentation.
 To further improve the accuracy of UNet-based models, researchers have developed several variants, such as attention UNet [18], UNet++ [1], and ResUNet++ [19].
@@
 [19] developed a ResUNet++ architecture, which takes advantage of residual blocks, squeeze and excitation blocks, atrous spatial pyramidal pooling, and attention blocks.
-Besides, several polyp segmentation methods [21, 22, 23, 24] have been developed and achieved promising performance.
+Moreover, the cross-level and multi-level feature fusion technique is widely recognized for enhancing segmentation performance by integrating spatial information from low-level features with semantic information from high-level features.
+For instance, Li et al.
+[20] proposed a cross-level information processing module that aggregates and processes multi-scale features transmitted by different encoder layers to improve polyp segmentation.
+Zhou et al.
+[21] designed a cross-level feature fusion module to merge adjacent features from various levels, enabling the characterization of cross-level and multi-scale information to address scale variations in polyps.
+Additionally, CFU-Net [22] incorporates a multi-level attention module to exploit interactions of contextual information, decision information, and long-range dependencies, thereby enhancing feature propagation within skip connections.
 II-B Semi-Supervised Learning
@@
 For example, Li et al.
-[27] proposed a segmentation model with multiple auxiliary decoders and encourage the consistency of the predictions made by the main decoder and the auxiliary decoders.
-Besides, contrastive learning [28, 17] has been widely used in SSL medical image segmentation.
+[25] proposed a segmentation model with multiple auxiliary decoders and encouraged the consistency of the predictions made by the main decoder and the auxiliary decoders.
+Besides, contrastive learning [26, 33, 34] has been widely used in SSL medical image segmentation.
 Moreover, adversarial learning [30, 31, 35] has also emerged to improve the robustness of the model.
@@
 Currently, several methods [10, 13, 12, 36, 37] based on the mean teacher framework focus on designing different consistent and perturbation strategies to achieve SSL segmentation.
-For instance, Verma et al.
-[36] regularized semi-supervised learning by encouraging consistent predictions at interpolations of unlabeled points u1 and u2.
+For instance, Tarvainen et al.
+[10] proposed to generate two different augmented views of the same input and use the output of the teacher model to supervise the student model.
+Chen et al.
+[13] proposed cross-pseudo supervision by enforcing the pseudo supervision between the predictions of two models with different parameter initialization.
 Wu et al.
@@
 [37] proposed a multi-attention tri-branch network (MTNet) that each branch uses a different attention mechanism.
-Among these methods, the consistent constraint is a widely used strategy, by enhancing that the perturbations on unlabeled data should not significantly vary their outputs.
+Among these methods, the consistent constraint is a widely used strategy, aiming to ensure that perturbations applied to unlabeled data do not drastically alter their outputs.
 […]
-Figure 2: The detailed architectures of the encoder (i.e., $E_{s}$ and $E_{t}$ in Fig.
-1) and scale-specific decoder (i.e., $D_{s1}$, $D_{s2}$, $D_{t1}$, and $D_{t2}$ in Fig.
+Figure 2: The detailed architectures of the encoder (i.e., $E$ in Fig.
+1) and scale-specific decoder (i.e., $D_{1}$ and $D_{2}$ in Fig.
 1) networks.
@@
 III-A Overview
-[…]
+The proposed DEC-Seg framework, as depicted in Fig.
+1, fully leverages multi-scale information from both labeled and unlabeled data for medical image segmentation.
+Our segmentation network comprises a shared encoder ($E$), two scale-specific decoders ($D_{1}$ processes features solely from the original scale while $D_{2}$ exclusively processes features from the downsampled scale), and a scale-fused decoder ($D_{f}$ is used to fuse features from both scales).
+Additionally, two generative networks ($G_{1}$ and $G_{2}$) assist in the segmentation process, enhancing the feature mining ability of our segmentation network through a cross-generation strategy.
+Specifically, the original and downsampled images from the labeled data are fed into the model (i.e., the encoder $E$ and scale-specific decoders $D_{1}$ and $D_{2}$), and then the scale-enhanced consistency and supervised loss are calculated to constrain the network.
+Then, the unlabeled data and its perturbed versions are inputted into the model.
+Meanwhile, we conduct scale-enhanced consistency, scale-aware perturbation consistency, and cross-generative consistency for unlabeled data, to fully leverage unlabeled samples to improve the segmentation performance.
+Moreover, to make use of multi-scale information, we design the scale-fused decoder (i.e., Df) to generate the final segmentation maps.
+For convenience, we denote ${X}_{l}=\left\{x_{l}^{1},x_{l}^{2},\cdots,x_{l}^{n_{l}}\right\}$ and ${Y}_{l}=\left\{y_{l}^{1},y_{l}^{2},\cdots,y_{l}^{n_{l}}\right\}$ as the labeled dataset and the corresponding label set, respectively, where $n_{l}$ is the number of the labeled images.
+We use ${X}_{u}=\left\{x_{u}^{1},x_{u}^{2},\cdots,x_{u}^{n_{u}}\right\}$ to denote the unlabeled dataset with $n_{u}$ images, and we typically have $n_{l}\ll n_{u}$.
+Given inputs $X$, they are fed into the encoder to learn five levels of features, namely $\{F_{i}\}_{i=1}^{5}$.
 III-B Cross-level Feature Aggregation Module
-To fully exploit cross-level information and provide more effective supplementary from different convolutional layers, we propose the CFA module to fuse the features from every two adjacent layers.
-Then, the aggregated feature can be incorporated into our designed decoder, and the feature flow is shown in Fig.
+To fully exploit cross-level information from various receptive fields and provide more effective supplementation from different convolutional layers, we propose the CFA module.
+This module fuses features from every two adjacent layers, enhancing the understanding of context and relationships within the cross-level features.
+The aggregated feature is then incorporated into our designed decoder, as depicted in Fig.
 2.
@@
 […]
-Figure 3: Illustrations of the encoder and cross-level feature aggregation module. “GAP” denotes the global average pooling, and “PWC” denotes the point-wise convolution.
+Figure 3: Illustrations of cross-level feature aggregation module.
+“GAP” denotes the global average pooling, “PWC” denotes the point-wise convolution, and “$U$” denotes the upsampling using bilinear interpolation.
 III-C Scale-enhanced Consistency
-The scale variation is still a great challenge for polyp segmentation.
+The scale variation is still a great challenge for medical image segmentation.
 As discussed in [39], reducing the gap between the network outputs from different scale images can help the model learn scale-correlated features.
 Motivated by this observation, we propose a scale-enhanced consistency scheme to constrain the outputs of different scale images to be closed, in which the different scale features can be refined by each other.
-Specifically, given unlabeled images $X_{u}$, $X_{u}$ and its downsampled versions $X_{u}^{d}=Down(X_{u})$ are fed to a weight-shared encoder to extract two sets of multi-level features, and then we can obtain the predicted maps $S_{u}$ and $S_{u}^{d}$, respectively.
-Specifically, the two segmentation maps can be obtained by $S_{u}=D_{s1}(E_{s}(X_{u}))$ and $S_{u}^{d}=D_{s2}(E_{s}(Down(X_{u})))$, where $E_{s}(\cdot)$ is the encoder in the student model, and $Down(\cdot)$ denotes a $1/2$ downsampling operation.
-$D_{s1}$ and $D_{s2}$ are scale-specific decoders in the student model, which are independent of each other to increase the perturbation.
-Similarly, for labeled images $X_{l}$, we can obtain the predicted segmentation maps $S_{l}$ and $S_{l}^{d}$ from $X_{l}$ and its downsampled versions $X_{l}^{d}$ by $S_{l}=D_{s1}(E_{s}(X_{l}))$ and $S_{l}^{d}=D_{s2}(E_{s}(Down(X_{l})))$, respectively.
-Therefore, we form a scale-enhanced consistency loss ($\mathcal{L}_{SC}$) to enforce $S$ and $S^{d}$ to be consistent, which is defined by
-$\mathcal{L}_{SC}=\mathcal{L}_{mse}(S_{u},S_{u}^{d})+\mathcal{L}_{mse}(S_{l},S_{l}^{d}),$ (3)
-where $\mathcal{L}_{mse}$ denotes the widely used mean square error loss. Finally, the supervised loss is formulated by
-$\begin{split}\mathcal{L}_{S}=&\mathcal{L}_{sup}(S_{l},Y_{l})+\mathcal{L}_{sup}(S_{l}^{d},Down(Y_{l})).\end{split}$ (4)
+To increase the perturbation, we use two independent decoders for the single-scale prediction task, which are called scale-special decoders (i.e., $D_{1}$ and $D_{2}$).
+Specifically, given unlabeled images $X_{u}$, $X_{u}$ and its downsampled versions $X_{u}^{2}=Down(X_{u})$ are fed to a weight-shared encoder to extract two sets of multi-level features, and then we can obtain the predicted maps $S_{u}^{1}$ and $S_{u}^{2}$, respectively.
+Specifically, the two segmentation maps can be obtained by $S_{u}^{1}=D_{1}(E(X_{u}))$ and $S_{u}^{2}=D_{2}(E(X_{u}^{2}))$, where $E(\cdot)$ is the encoder, and $Down(\cdot)$ denotes a $1/2$ downsampling operation.
+Furthermore, we employ the mutual supervision approach outlined in [13] to ensure the consistency of $S^{1}_{u}$ and $S^{2}_{u}$ with different sizes.
+Concretely, $\hat{Y}^{1}_{u}=\arg\max\limits_{c}(S^{1}_{u})$ and $\hat{Y}^{2}_{u}=\arg\max\limits_{c}(S^{2}_{u})$ are computed.
+Subsequently, $\hat{Y}^{1}_{u}$ is downsampled to the size of $S^{2}_{u}$ and utilized as a pseudo-label for $S^{2}_{u}$, while $\hat{Y}^{2}_{u}$ is upsampled to the size of $S^{1}_{u}$ and employed as a pseudo-label for $S^{1}_{u}$.
+Therefore, the scale-enhanced consistency loss for unlabeled data is defined by
+$\mathcal{L}_{SC}^{u}(S^{1}_{u},S^{2}_{u})=\mathcal{L}_{ce}(S^{1}_{u},Up(\hat{Y}^{2}_{u}))+\mathcal{L}_{ce}(S^{2}_{u},Down(\hat{Y}^{1}_{u})),$ (3)
+where $\mathcal{L}_{ce}$ denotes the widely used cross entropy loss and $Up(\cdot)$ denotes a 2×2$\times$2 × upsampling operation.
+Similarly, for labeled images $X_{l}$, we can obtain the predicted segmentation maps $S_{l}^{1}$ and $S_{l}^{2}$ from $X_{l}$ and its downsampled versions $X_{l}^{2}$ by $S_{l}^{1}=D_{1}(E(X_{l}))$ and $S_{l}^{2}=D_{2}(E(X_{l}^{2}))$, respectively.
+As a result, $\mathcal{L}_{SC}^{l}(S^{1}_{l},S^{2}_{l})$ can be obtained in the same way.
+[…]
+Figure 4: Illustration of (a) the architecture of the scale-fused decoder (i.e., ${D}_{f}$ in Fig.
+1) and (b) the proposed dual-scale complementary fusion module.
+“$U$” denotes the upsampling using a bilinear interpolation.
 III-D Dual-scale Complementary Fusion Module
-[…]
+The two independent decoders, as designed in Section III-C, are specialized for processing information at two distinct scales.
+Throughout network training, these decoders acquire unique features based on the input image scale, with each emphasizing different aspects.
+Despite this distinction, the features from both scales pertain to the same unlabeled image, indicating the opportunity to combine and extract features from these two scale features further, potentially resulting in an enhanced segmentation map.
+Consequently, to fully leverage information from different scales, in comparison to the scale-specific decoder, we also introduce the scale-fused decoder (Df in Fig.
+1) to produce more reliable predictions as the final segmentation maps.
+As shown in Fig.
+1, the features from $D_{1}$ and $D_{2}$ are fused and then fed into the scale-fused decoder.
+The feature flow of the scale-fused decoder is shown in Fig.
+4(a), and the structure of the scale-fused decoder is basically the same as that of the scale-specific decoder.
+To achieve the fuse operation, we present a dual-scale complementary fusion (DCF) module to fuse the features from the two scale-specific decoders.
+This module replaces the features transmitted from the encoder through the skip operation.
+The proposed DCF module aims to fuse the multi-scale features and enable the features from the original scale and downsampled scale to complement each other (as shown in Fig.
+4(b)).
+Taking the original scale feature as an example, we first conduct a $3\times 3$ convolution on $Up(h_{i}^{2})$ and Sigmoid activation function to obtain a scale-aware weight $W_{2}=\sigma(\mathcal{B}_{conv3\times 3}(Up(h_{i}^{2})))$.
+Then, we multiply $W_{2}$ by $\mathcal{B}_{conv3\times 3}(Up(h_{i}^{2}))$, and the resultant feature is the complementary information (i.e., $h_{com}^{2}$) which is encouraged to learn the information required by the original scale.
+Finally, we add the complementary feature obtained at the downsampled scale to the original scale and then smooth the feature by a $3\times 3$ convolution, thus we can obtain $h_{rec}^{1}=\mathcal{B}_{conv3\times 3}(\mathcal{B}_{conv3\times 3}(h_{i}^{1})+\mathcal{B}_{conv3\times 3}(Up(h_{i}^{2}))*W_{2})$.
+Similarly, we can obtain the complementary-enhanced downsampled feature $h_{rec}^{2}$.
+Through a cross-scale complementary-enhanced process, we learn more rich multi-scale feature representations.
+Further, two features $h_{rec}^{1}$ and $h_{rec}^{2}$ are concatenated and then proceeded by a $1\times 1$ convolution to two channels and a Softmax function to obtain two weight maps $\alpha_{1}$ and $\alpha_{2}$, where $\alpha_{1}+\alpha_{2}=1$.
+Finally, we obtain the multi-scale fused feature by
 $h_{i}^{dcf}=h_{rec}^{1}\otimes\alpha_{1}+h_{rec}^{2}\otimes\alpha_{2}.$ (4)
-Subsequently, the fused features $h_{i}^{dcf}$ will be passed into the fused decoder.
-Specifically, $h_{5}^{dcf}$ is first passed through a transpose convolutional layer to have the same resolution with $h_{4}^{dcf}$.
-Then, the two features are concatenated and fed into two convolutional layers.
-Repeating the above process, we will obtain the final segmentation maps $S_{l}^{f}$ for labeled data.
-Therefore, the supervised loss in Eq.
-(4) can be reformulated as follows:
-$\begin{split}\mathcal{L}_{S}=&\mathcal{L}_{sup}(S_{l},Y_{l})+\\ &\mathcal{L}_{sup}(S_{l}^{d},Down(Y_{l}))+\mathcal{L}_{sup}(S_{l}^{f},Y_{l}).\end{split}$ (6)
-[…]
-Figure 4: Illustration of (a) the architecture of the fused decoder (i.e., ${D}_{s}^{f}$ and ${D}_{t}^{f}$ in Fig. 1) and (b) the proposed dual-scale complementary fusion module.
+Subsequently, the fused features $h_{i}^{dcf}$ will concatenate with the features from the previous layer and continue to propagate forward.
+Specifically, $h_{5}^{dcf}$ is first passed through two convolutional layers and a transpose convolutional layer to have the same resolution with $h_{4}^{dcf}$.
+Then, the two features are concatenated and fed into the next two convolutional layers.
+Repeating the above process, we will obtain the final segmentation maps $S^{f}$.
+It is important to highlight that the scale-fused decoder plays a vital role in fully integrating multi-scale information, consequently enhancing the segmentation performance.
+Consequently, during the inference stage, the segmentation predictions obtained from the scale-fused decoder are taken as the final segmentation results.
 […]
@@
 III-E Scale-aware Perturbation and Cross-generative Consistency
-Following previous works [35, 33, 27, 42, 43], to improve the robustness of the MT model, $X_{u}$ and its perturbed version $X_{u}^{p}$ are passed through both student and teacher models, respectively, to encourage the outputs to be closed.
-In this study, we design a scale-aware perturbation consistency.
-Specifically, we impose the perturbation consistency on different scale images, i.e., $S_{u}^{p}=D_{t1}(E_{t}(\mathcal{P}(X_{u})))$ and $S_{u}^{dp}=D_{t2}(E_{t}(\mathcal{P}(Down(X_{u}))))$, where a random gamma operation is adopted as the perturbation operation (i.e., $\mathcal{P}$).
-Meanwhile, we also obtain ${S}_{u}^{fp}$ that its constraint is implemented by the MT framework, due to the particularity of ${S}_{u}^{fp}$.
-Thus, the scale-aware perturbation consistency scheme can be optimized by minimizing $\mathcal{L}_{SPC}$, which is defined by
-$\mathcal{L}_{SPC}=\mathcal{L}_{mse}(S_{u},S_{u}^{p})+\mathcal{L}_{mse}(S_{u}^{d},S_{u}^{dp})+\mathcal{L}_{mse}(S_{u}^{f},S_{u}^{fp}).$ (7)
-Further, to utilize the rich information contained in different feature maps, we propose a perturbation-based cross-generative consistency constraint to enhance the feature mining ability of our segmentation network.
+To constrain the predictions of unlabeled data on the scale-fused decoder, we follow previous works [25, 40, 41], and introduce scale-aware perturbation consistency to encourage the outputs between the original version and perturbed version to be closed.
+Specifically, we impose the perturbation consistency on different scale images, i.e., $S_{p}^{1}=D_{1}(E(\mathcal{P}(X_{u})))$, $S_{p}^{2}=D_{2}(E(\mathcal{P}(X_{u}^{2})))$, and $S_{p}^{f}$, where a random color jitter operation is adopted as the perturbation operation (i.e., $\mathcal{P}$).
+Thus, $\mathcal{L}_{SC}^{p}(S^{1}_{p},S^{2}_{p})$ is obtained and the scale-aware perturbation consistency scheme can be optimized by minimizing $\mathcal{L}_{SPC}$, which is defined by
+$\mathcal{L}_{SPC}=\mathcal{L}_{mse}(S_{u}^{1},S_{p}^{1})+\mathcal{L}_{mse}(S_{u}^{2},S_{p}^{2})+\mathcal{L}_{mse}(S_{u}^{f},S_{p}^{f}).$ (5)
+Further, to utilize the rich information contained in different feature maps, we propose a perturbation-based cross-generative consistency constraint to enhance the feature mining ability of our segmentation network with information reconstruction and to correct the cognitive bias on a single input version with cross-generation.
 Therefore, we design two generative networks $G_{1}$ and $G_{2}$ (The detailed structure is shown in Fig.
 5) for image reconstruction.
-The segmentation maps $S_{u}^{f}$ and $S_{u}^{fp}$ obtained from the aggregation of the two scale information are fed into $G_{2}$ and $G_{1}$, respectively, and then we can obtain two reconstructed images $\hat{X}_{u}=G_{1}(S_{u}^{fp})$ and $\hat{X}_{u}^{p}=G_{2}(S_{u}^{f})$.
+For the generative network to accept richer information, produce better-reconstructed maps, and deliver more valuable feedback to the segmentation network, the logit predictions $z_{u}^{f}$ and $z_{p}^{f}$ obtained from the scale-fused decoder are fed into $G_{2}$ and $G_{1}$, respectively, and then we can obtain two reconstructed images $\hat{X}_{p}=G_{1}(z_{p}^{f})$ and $\hat{X}_{u}=G_{2}(z_{u}^{f})$.
 Finally, we form a cross-generative consistency loss ($\mathcal{L}_{CC}$), which is given by
-$\mathcal{L}_{CC}=\mathcal{L}_{mse}(\hat{X}_{u},X_{u})+\mathcal{L}_{mse}(\hat{X}_{u}^{p},X_{u}^{p}).$ (8)
-It is worth noting that $\mathcal{L}_{CC}$ can enforce $\hat{X}_{u}$ and $X_{u}$ to be consistent and $\hat{X}_{u}^{p}$ and $X_{u}^{p}$ to be closed.
-As a result, such cross-generation can further improve the robustness of our MT-based model to learn powerful features and fully harness the knowledge from unlabeled data to boost the proposed segmentation model.
+$\mathcal{L}_{CC}=\mathcal{L}_{mse}(\hat{X}_{p},X_{u})+\mathcal{L}_{mse}(\hat{X}_{u},X_{p}).$ (6)
+It is worth noting that $\mathcal{L}_{CC}$ can enforce $\hat{X}_{p}$ and $X_{u}$ to be consistent and $\hat{X}_{u}$ and $X_{p}$ to be closed.
+Using the input image as a learning target provides unlabeled data with an authentic constraint containing structural and semantic information.
+This empowers the segmentation network to prioritize the structural and semantic details that are pertinent to category information during the generation of segmentation maps.
+The cross-generation process compels the network to investigate inconsistent predictions caused by color variations, thereby rectifying the network’s cognitive biases toward a single input and ensuring the accuracy of predictions for both the original input and the perturbed input from different perspectives.
+Consequently, such cross-generation can further enhance the robustness of our model, enabling it to learn powerful features and effectively leverage the knowledge from unlabeled data to enhance the proposed segmentation model.
+III-F Overall Loss Function
+In this study, the supervised loss for labeled data is the sum of the loss on the prediction maps of the designed three decoders and its corresponding ground truth, and it is formulated by
+$\begin{split}\mathcal{L}_{S}=\mathcal{L}_{sup}(S_{l}^{1},Y_{l})+\mathcal{L}_{sup}(S_{l}^{2},Down(Y_{l}))+\mathcal{L}_{sup}(S_{l}^{f},Y_{l}),\end{split}$ (7)
+where $\mathcal{L}_{sup}=(\mathcal{L}_{CE}+\mathcal{L}_{Dice})/2$, $\mathcal{L}_{CE}$ and $\mathcal{L}_{Dice}$ denote the cross-entropy (CE) loss and Dice loss, respectively.
+Finally, the total loss can be given as follows:
+$\mathcal{L}_{total}=\mathcal{L}_{S}+\mathcal{L}_{SPC}+\mathcal{L}_{SC}+\mathcal{L}_{CC},$ (8)
+where $\mathcal{L}_{SC}=\mathcal{L}_{SC}^{l}+\mathcal{L}_{SC}^{u}+\mathcal{L}_{SC}^{p}$ is the sum of scale-enhanced consistency loss in labeled images, unlabeled images, and perturbed images.
+IV Experiments and Results
+IV-A Datasets
+Colonoscopy Datasets: We carry out the comparison experiments on four public datasets, i.e., CVC-ColonDB [42], Kvasir [43], CVC-ClinicDB [44], and ETIS [45].
+Following the setting in [46], $1,450$ images are randomly selected from the two datasets (i.e., CVC-ClinicDB and Kvasir) for the training set, and the remaining images from the two datasets and the other two datasets (i.e., ETIS and CVC-ColonDB) to form the testing set.
+Besides, $10\%$ ($145$ images) or $30\%$ ($435$ images) from the training set are used as the labeled data, while the remaining images are adopted as the unlabeled data.
+Brain Tumor Dataset: This dataset contains brain MR images together with manual FLAIR abnormality segmentation masks and it comes from 110 patients included in The Cancer Genome Atlas (TCGA) lower-grade glioma [47, 48].
+In this study, we have $1,090$ images for the training set and 283 ones for the testing set.
+In addition, $10\%$ and $30\%$ images from the training set are adopted as the labeled data, respectively, while the remaining training images are taken as unlabeled data.
+Skin Lesion Dataset: The dataset comes from ISIC-2018 challenge [49] and is used for skin lesion segmentation.
+It includes $2,594$ dermoscopic images for training and $1,000$ images for testing.
+In the same way, $10\%$ or $30\%$ images from the training set are used as the labeled data, while the remaining images are formed as the unlabeled set.
 TABLE I: Quantitative results on CVC-ClinicDB and Kvasir datasets with $10\%$ or $30\%$ labeled data.
@@
 mDice $\uparrow$ mIoU $\uparrow$ $F_{\beta}^{w}$uparrow $S_{\alpha}$uparrow MAE $\downarrow$ mDice $\uparrow$ mIoU $\uparrow$ $F_{\beta}^{w}$uparrow $S_{\alpha}$uparrow MAE $\downarrow$
-CVC-ClinicDB Baseline 0.767 0.689 0.756 0.847 0.024 0.812 0.751 0.806 0.873 0.021
-MT [15] 0.785 0.711 0.763 0.855 0.025 0.839 0.781 0.829 0.892 0.020
-DAN [34] 0.789 0.718 0.779 0.862 0.024 0.838 0.781 0.835 0.890 0.018
-UA-MT [10] 0.798 0.729 0.782 0.869 0.022 0.831 0.762 0.817 0.879 0.026
-URPC [16] 0.794 0.723 0.783 0.864 0.022 0.829 0.767 0.825 0.880 0.022
-Duo-SegNet [32] 0.765 0.685 0.745 0.840 0.037 0.755 0.678 0.739 0.832 0.038
+CVC-ClinicDB MT [10] 0.747 0.662 0.741 0.830 0.028 0.772 0.689 0.763 0.852 0.028
+DAN [32] 0.755 0.674 0.733 0.842 0.027 0.775 0.696 0.772 0.848 0.027
+UA-MT [11] 0.749 0.676 0.742 0.839 0.027 0.790 0.715 0.787 0.853 0.025
+URPC [14] 0.769 0.696 0.760 0.854 0.021 0.777 0.709 0.785 0.853 0.026
 CLCC [15] 0.794 0.720 0.786 0.859 0.027 0.840 0.786 0.840 0.891 0.019
@@
 CDMA [37] 0.759 0.676 0.750 0.837 0.025 0.803 0.728 0.802 0.863 0.024
-BCP [44] 0.835 0.769 0.828 0.887 0.018 0.846 0.782 0.847 0.894 0.020
-DEC-Seg (ours) 0.830 0.772 0.836 0.888 0.014 0.852 0.798 0.853 0.897 0.019
-Kvasir Baseline 0.786 0.701 0.756 0.831 0.059 0.843 0.782 0.829 0.877 0.037
-MT [15] 0.833 0.755 0.818 0.858 0.051 0.868 0.799 0.853 0.881 0.037
-DAN [34] 0.836 0.758 0.820 0.858 0.049 0.867 0.803 0.851 0.886 0.035
-UA-MT [10] 0.824 0.753 0.808 0.859 0.051 0.858 0.787 0.838 0.881 0.039
-URPC [16] 0.821 0.740 0.803 0.853 0.050 0.854 0.789 0.843 0.879 0.037
-Duo-SegNet [32] 0.795 0.711 0.785 0.838 0.061 0.851 0.772 0.843 0.878 0.043
+SCP-Net [50] 0.776 0.703 0.758 0.854 0.024 0.839 0.786 0.837 0.888 0.020
+MCF [51] 0.779 0.718 0.767 0.860 0.021 0.823 0.766 0.819 0.882 0.019
+DEC-Seg (ours) 0.836 0.774 0.834 0.886 0.019 0.859 0.804 0.860 0.900 0.015
+Kvasir MT [10] 0.814 0.722 0.797 0.841 0.050 0.814 0.732 0.790 0.852 0.052
+DAN [32] 0.808 0.723 0.786 0.839 0.059 0.841 0.760 0.828 0.866 0.043
+UA-MT [11] 0.799 0.713 0.782 0.834 0.058 0.845 0.771 0.840 0.873 0.037
+URPC [14] 0.811 0.728 0.796 0.842 0.057 0.849 0.778 0.846 0.874 0.040
 CLCC [15] 0.806 0.724 0.796 0.844 0.052 0.864 0.804 0.853 0.892 0.034
@@
 CDMA [37] 0.786 0.695 0.769 0.826 0.056 0.839 0.758 0.826 0.862 0.044
-BCP [44] 0.854 0.789 0.850 0.876 0.042 0.886 0.820 0.876 0.898 0.029
-DEC-Seg (ours) 0.862 0.794 0.853 0.880 0.040 0.882 0.822 0.871 0.899 0.033
-III-F Loss Function and Optimization
-Inspired by [45], the supervised loss for labeled data is formulated by $\mathcal{L}_{sup}=\mathcal{L}_{IoU}^{w}+\mathcal{L}_{BCE}^{w}$, where $\mathcal{L}_{IoU}^{w}$ and $\mathcal{L}_{BCE}^{w}$ denote the weighted IoU loss and binary cross-entropy (BCE) loss, respectively.
-Finally, the total loss can be given as follows:
-$\mathcal{L}_{total}=\mathcal{L}_{S}+\mathcal{L}_{SPC}+\alpha\mathcal{L}_{SC}+\beta\mathcal{L}_{CC},$ (9)
-where the hyper-parameter $\alpha$ is set to $1.5$ and $\beta$ is set to $3$.
-In addition, the parameters $\theta^{{}^{\prime}}$ of the teacher network can be updated based on the parameters $\theta$ of the student network using an exponential moving average (EMA) strategy.
-Thus, at the $t$-th training step, the parameters $\theta^{{}^{\prime}}$ can be updated by $\theta^{{}^{\prime}}_{t}=\gamma\theta^{{}^{\prime}}_{t-1}+(1-\gamma)\theta_{t}$, where $\gamma$ is used to control the EMA decay, and it is set to $0.99$.
-IV EXPERIMENTS AND RESULTS
-IV-A Experimental Setup
-IV-A1 Datasets and Training Settings
-We carry out the comparison experiments on five public datasets, i.e., ETIS [46], CVC-ClinicDB [47], EndoScene-CVC300 [48], CVC-ColonDB [49], and Kvasir [50].
-Following the setting in [2], $1450$ images are randomly selected from the two datasets (Kvasir and CVC-ClinicDB) for the training set, and the remaining images from the two datasets and the other three datasets (EndoScene-CVC300, ETIS, and CVC-ColonDB) to form the testing set.
-Besides, $10\%$ ($145$ images) or $30\%$ ($435$ images) from the training set are used as the labeled data, while the remaining images are adopted as the unlabeled data.
-IV-A2 Implementation Details and Evaluation Metrics
-The proposed segmentation framework is implemented in PyTorch and trained on one NVIDIA GeForce RTX3090 GPU with an Adam optimizer.
-All inputs are uniformly rescaled to $352\times 352$.
-Our model converges over $10000$ iterations with a batch size of $8$ including 4 labeled samples and 4 unlabeled samples, and the learning rate is set to $1e$-$4$.
-To evaluate the effectiveness, we employ five commonly adopted metrics [2, 3], namely mean Dice (mDice), mean IoU (mIoU), $F_{\beta}^{w}$, $S_{\alpha}$, and mean absolute error (MAE).
-TABLE II: Quantitative results on three unseen datasets (EndoScene-CVC300, CVC-ColonDB, and ETIS) with $10\%$ or $30\%$ labeled data.
+SCP-Net [50] 0.810 0.723 0.790 0.840 0.056 0.862 0.789 0.847 0.878 0.037
+MCF [51] 0.822 0.751 0.816 0.852 0.051 0.856 0.794 0.847 0.882 0.039
+DEC-Seg (ours) 0.859 0.787 0.853 0.877 0.044 0.893 0.830 0.886 0.902 0.032
+IV-B Implementation Details and Evaluation Metrics
+Implementation Details: The proposed segmentation framework is implemented in PyTorch and trained on one NVIDIA GeForce RTX3090 GPU.
+The input images from the polyp segmentation dataset and skin lesion segmentation dataset are uniformly rescaled to $352\times 352$, and the inputs of the brain MRI segmentation dataset are uniformly rescaled to $256\times 256$.
+Our model converges over $10,000$ iterations with a batch size of $6$ including 3 labeled samples and 3 unlabeled samples, and uses an SGD optimizer with an initial learning rate set to $1e$-$2$ and a poly learning rate strategy to update the learning rate.
+Evaluation Metrics: To evaluate the effectiveness, we employ five commonly adopted metrics [46, 19], namely mean Dice (mDice), mean IoU (mIoU), $F_{\beta}^{w}$, $S_{\alpha}$, and mean absolute error (MAE).
+TABLE II: Quantitative results on two unseen datasets (CVC-ColonDB and ETIS) with $10\%$ or $30\%$ labeled data.
 Methods 10% labeled 30% labeled
 mDice $\uparrow$ mIoU $\uparrow$ $F_{\beta}^{w}$uparrow $S_{\alpha}$uparrow MAE $\downarrow$ mDice $\uparrow$ mIoU $\uparrow$ $F_{\beta}^{w}$uparrow $S_{\alpha}$uparrow MAE $\downarrow$
-EndoScene-CVC300 Baseline 0.800 0.708 0.784 0.861 0.013 0.827 0.760 0.812 0.895 0.008
-MT [15] 0.858 0.779 0.837 0.910 0.010 0.873 0.796 0.845 0.916 0.009
-DAN [34] 0.820 0.733 0.797 0.885 0.014 0.874 0.800 0.847 0.917 0.008
-UA-MT [10] 0.859 0.780 0.829 0.910 0.010 0.837 0.754 0.795 0.895 0.012
-URPC [16] 0.841 0.753 0.812 0.896 0.012 0.872 0.800 0.859 0.920 0.007
-Duo-SegNet [32] 0.841 0.761 0.811 0.894 0.011 0.843 0.758 0.803 0.902 0.012
-CLCC [17] 0.748 0.674 0.724 0.845 0.015 0.826 0.755 0.811 0.891 0.008
-SLC-Net [18] 0.856 0.777 0.827 0.907 0.012 0.869 0.796 0.849 0.915 0.010
-MC-Net+ [37] 0.783 0.711 0.755 0.872 0.014 0.845 0.783 0.844 0.891 0.020
-CDMA [38] 0.791 0.685 0.765 0.846 0.013 0.803 0.728 0.802 0.863 0.024
-BCP [44] 0.856 0.770 0.835 0.901 0.010 0.846 0.782 0.847 0.894 0.020
-DEC-Seg (ours) 0.860 0.781 0.831 0.914 0.011 0.895 0.824 0.875 0.927 0.006
-CVC-ColonDB Baseline 0.531 0.453 0.526 0.712 0.052 0.551 0.489 0.543 0.732 0.049
-MT [15] 0.610 0.531 0.597 0.759 0.044 0.668 0.587 0.652 0.788 0.041
-DAN [34] 0.599 0.515 0.593 0.748 0.049 0.629 0.551 0.613 0.766 0.043
-UA-MT [10] 0.583 0.507 0.576 0.745 0.050 0.666 0.587 0.641 0.785 0.048
-URPC [16] 0.587 0.507 0.574 0.745 0.048 0.621 0.551 0.612 0.768 0.045
-Duo-SegNet [32] 0.644 0.550 0.613 0.763 0.058 0.658 0.569 0.627 0.773 0.057
+CVC-ColonDB MT [10] 0.589 0.497 0.581 0.743 0.048 0.600 0.510 0.597 0.749 0.045
+DAN [32] 0.620 0.517 0.604 0.756 0.046 0.633 0.546 0.630 0.767 0.045
+UA-MT [11] 0.553 0.471 0.551 0.727 0.051 0.648 0.563 0.635 0.771 0.046
+URPC [14] 0.556 0.480 0.549 0.732 0.047 0.598 0.519 0.601 0.754 0.045
 CLCC [15] 0.538 0.473 0.533 0.720 0.053 0.564 0.505 0.562 0.738 0.050
@@
 CDMA [37] 0.507 0.419 0.494 0.691 0.054 0.624 0.536 0.609 0.756 0.047
-BCP [44] 0.642 0.563 0.635 0.773 0.046 0.656 0.580 0.651 0.778 0.046
-DEC-Seg (ours) 0.650 0.575 0.639 0.782 0.042 0.691 0.616 0.678 0.798 0.042
-ETIS Baseline 0.458 0.392 0.433 0.693 0.024 0.486 0.432 0.469 0.715 0.021
-MT [15] 0.468 0.395 0.434 0.694 0.036 0.595 0.519 0.552 0.761 0.035
-DAN [34] 0.473 0.395 0.440 0.700 0.031 0.528 0.452 0.485 0.729 0.045
-UA-MT [10] 0.469 0.399 0.444 0.695 0.027 0.603 0.515 0.547 0.761 0.032
-URPC [16] 0.459 0.391 0.436 0.691 0.028 0.601 0.523 0.572 0.765 0.023
-Duo-SegNet [32] 0.450 0.382 0.407 0.669 0.061 0.578 0.500 0.531 0.750 0.049
+SCP-Net [50] 0.577 0.495 0.559 0.739 0.049 0.694 0.615 0.683 0.802 0.039
+MCF [51] 0.566 0.498 0.559 0.742 0.045 0.621 0.550 0.614 0.768 0.042
+DEC-Seg (ours) 0.648 0.565 0.630 0.771 0.046 0.721 0.640 0.709 0.814 0.035
+ETIS MT [10] 0.356 0.288 0.337 0.636 0.035 0.495 0.420 0.462 0.706 0.039
+DAN [32] 0.437 0.358 0.397 0.666 0.046 0.485 0.418 0.472 0.708 0.027
+UA-MT [11] 0.459 0.393 0.441 0.693 0.029 0.541 0.473 0.523 0.738 0.024
+URPC [14] 0.420 0.356 0.406 0.677 0.028 0.559 0.490 0.549 0.749 0.021
 CLCC [15] 0.409 0.346 0.395 0.668 0.031 0.474 0.422 0.465 0.707 0.020
@@
 CDMA [37] 0.309 0.254 0.296 0.611 0.031 0.531 0.446 0.494 0.724 0.024
-BCP [44] 0.571 0.502 0.550 0.756 0.023 0.557 0.497 0.541 0.754 0.020
-DEC-Seg (ours) 0.615 0.540 0.582 0.775 0.019 0.683 0.609 0.658 0.814 0.014
-IV-B Comparison with State-of-the-arts
-We compare the proposed DEC-Seg with ten state-of-the-art semi-supervised segmentation methods, i.e., MT [15], DAN [34], UA-MT [10], URPC [16], Duo-SegNet [32], CLCC [17], SLC-Net [18], MC-Net+ [37], CDMA [38] and BCP [44].
+SCP-Net [50] 0.394 0.323 0.355 0.656 0.045 0.571 0.500 0.543 0.747 0.023
+MCF [51] 0.425 0.367 0.407 0.679 0.025 0.487 0.425 0.469 0.707 0.022
+DEC-Seg (ours) 0.592 0.511 0.558 0.758 0.025 0.634 0.564 0.608 0.789 0.019
+TABLE III: Quantitative results on Brain MRI dataset with $10\%$ or $30\%$ labeled data.
+Methods 10% labeled 30% labeled
+mDice $\uparrow$ mIoU $\uparrow$ $F_{\beta}^{w}$uparrow $S_{\alpha}$uparrow MAE $\downarrow$ mDice $\uparrow$ mIoU $\uparrow$ $F_{\beta}^{w}$uparrow $S_{\alpha}$uparrow MAE $\downarrow$
+MT [10] 0.491 0.404 0.512 0.705 0.015 0.674 0.561 0.704 0.792 0.012
+DAN [32] 0.534 0.429 0.564 0.719 0.015 0.703 0.582 0.737 0.803 0.011
+UA-MT [11] 0.549 0.435 0.584 0.721 0.015 0.703 0.586 0.736 0.807 0.011
+URPC [14] 0.547 0.448 0.571 0.729 0.014 0.653 0.548 0.677 0.785 0.012
+CLCC [15] 0.595 0.505 0.615 0.760 0.012 0.728 0.615 0.754 0.824 0.010
+SLC-Net [16] 0.483 0.400 0.506 0.704 0.015 0.702 0.591 0.714 0.807 0.010
+MC-Net+ [36] 0.580 0.481 0.595 0.749 0.015 0.746 0.640 0.766 0.834 0.009
+CDMA [37] 0.495 0.408 0.508 0.705 0.015 0.627 0.526 0.647 0.775 0.011
+SCP-Net [50] 0.559 0.467 0.571 0.741 0.014 0.643 0.544 0.660 0.783 0.012
+MCF [51] 0.589 0.491 0.623 0.750 0.013 0.704 0.612 0.722 0.819 0.010
+DEC-Seg (ours) 0.626 0.533 0.639 0.777 0.012 0.767 0.671 0.783 0.850 0.008
+TABLE IV: Quantitative results on ISIC-2018 dataset with $10\%$ or $30\%$ labeled data.
+Methods 10% labeled 30% labeled
+mDice $\uparrow$ mIoU $\uparrow$ $F_{\beta}^{w}$uparrow $S_{\alpha}$uparrow MAE $\downarrow$ mDice $\uparrow$ mIoU $\uparrow$ $F_{\beta}^{w}$uparrow $S_{\alpha}$uparrow MAE $\downarrow$
+MT [10] 0.822 0.738 0.809 0.811 0.104 0.844 0.760 0.822 0.825 0.087
+DAN [32] 0.831 0.744 0.817 0.815 0.097 0.842 0.757 0.825 0.824 0.091
+UA-MT [11] 0.839 0.752 0.818 0.820 0.094 0.845 0.762 0.829 0.829 0.086
+URPC [14] 0.847 0.763 0.829 0.831 0.089 0.853 0.772 0.835 0.880 0.082
+CLCC [15] 0.842 0.756 0.815 0.820 0.088 0.845 0.762 0.816 0.821 0.092
+SLC-Net [16] 0.843 0.754 0.818 0.822 0.091 0.846 0.752 0.811 0.821 0.085
+MC-Net+ [36] 0.848 0.767 0.829 0.832 0.087 0.862 0.782 0.841 0.842 0.076
+CDMA [37] 0.858 0.777 0.842 0.839 0.083 0.868 0.788 0.848 0.846 0.073
+SCP-Net [50] 0.852 0.770 0.835 0.835 0.085 0.853 0.769 0.831 0.834 0.083
+MCF [51] 0.856 0.773 0.837 0.836 0.083 0.861 0.778 0.839 0.840 0.077
+DEC-Seg (ours) 0.867 0.796 0.858 0.851 0.078 0.875 0.801 0.862 0.856 0.069
+IV-C Comparison with State-of-the-arts
+We compare the proposed DEC-Seg with ten state-of-the-art semi-supervised segmentation methods, i.e., MT [10], DAN [32], UA-MT [11], URPC [14], CLCC [15], SLC-Net [16], MC-Net+ [36], CDMA [37], SCP-Net [50] and MCF [51].
 For a fair comparison, we change the backbones of all compared methods to “Res2Net”, and then train all compared methods without any data augmentation.
-Moreover, we adopt a basic U-Net [19] model with Res2Net as the backbone of the encoder, denoted as “Baseline”.
-IV-B1 Quantitative Results
+All the compared methods undergo $10,000$ iterations, similar to our approach, utilizing the same data and image dimensions.
+In the case of CLCC, we maintain its original setting of $320\times 320$ due to the specificity of the method.
+Moreover, the optimizer, learning rate, and descent strategy all remain consistent with their original configurations across the compared methods.
+IV-C1 Experiments on Polyp Segmentation
 Quantitative comparisons are reported in Tab.
@@
 II.
-To validate the learning ability of the two seen training datasets (CVC-ClinicDB and Kvasir), we can be seen from Tab.
-I that while most methods learn well, our method performs better than other compared methods, both on $10\%$ and $30\%$ of labeled data.
-In addition, due to the additional unlabeled data, our method improves mDice and mIoU by $8.2\%$ and $12.0\%$ compared to baseline using only $10\%$ labeled data, and by $4.9\%$ and $6.3\%$ compared to baseline using only $30\%$ labeled data on CVC-ClinicDB, which indicate that our DEC-Seg can fully leverage unlabeled data to improve the polyp segmentation performance.
-Additionally, we report the comparison results on three unseen datasets (CVC-300, ETIS, and CVC-ColonDB) from Tab.
+To validate the learning ability of the two seen training datasets (CVC-ClinicDB and Kvasir), it can be seen from Tab.
+I that while most methods learn well, our method performs better than other compared methods, both on $10\%$ and $30\%$ of labeled data, which indicate that our DEC-Seg can fully leverage unlabeled data to improve the polyp segmentation performance.
+Additionally, we report the comparison results on two unseen datasets (ETIS, and CVC-ColonDB) from Tab.
 II to verify the generalization ability of our DEC-Seg.
-Because the polyp in the CVC-300 dataset is similar to the training set and relatively simple, the metrics are also similar to CVC-ClinicDB and Kvisir datasets.
-Compared with SLC-Net, which performs relatively stable on this dataset, our method achieves $3.0\%$ and $3.5\%$ improvements in terms of mDice and mIoU with $30\%$ labeled data.
-The CVC-ColonDB and ETIS datasets are more difficult, as polyps are with large variations in size.
+The CVC-ColonDB and ETIS datasets are difficult, as polyps are with large variations in size.
 As reported in Tab.
-II, comparing the BCP method on the CVC-ColonDB dataset using $10\%$ labeled data, our method achieves $1.2\%$ and $2.1\%$ improvements on mDice and mIoU, respectively.
-Besides, our method achieves $5.3\%$ and $6.2\%$ improvements over the SOTA semi-supervised segmentation method (BCP) in terms of mDice and mIoU with the $30\%$ labeled data.
+II, comparing the SCP-Net method on the CVC-ColonDB dataset using $30\%$ labeled data, our method achieves $3.9\%$ and $4.1\%$ improvements on mDice and mIoU, respectively.
 Similarly, on the ETIS dataset, our performance has a more significant superiority over other methods.
-The main reason is that our exploration and utilization of scale information make the model more robust to scale variation, which is able to identify small polyps that cannot be identified by other methods, and can also refine the edges of large polyps.
+The primary reason is that our exploration and utilization of scale information enhance the model’s robustness to scale variations.
+This capability allows it to small polyps, while other methods cannot identify them accurately.
 Overall, the results indicate that our DEC-Seg has a better generalization ability and effectively utilizes unlabeled data to boost polyp segmentation.
 […]
-Figure 6: Qualitative results of our model and other ten compared semi-supervised segmentation methods.
-IV-B2 Qualitative Results
+Figure 6: Qualitative results of our model and other ten compared semi-supervised segmentation methods on polyp segmentation datasets.
 Fig.
-6 shows some of the visualization results of the test examples under $30\%$ labeled data.
+6 presents some of the visualization results of the test examples under $30\%$ labeled data.
 It can be seen that our method can accurately locate and segment polyps under different challenging factors.
-For example, in the first two rows of Fig.
+For example, in the first row of Fig.
 6, the polyps have very small sizes.
-It can be clearly seen that Duo-SegNet recognizes the polyp incorrectly, while UAMT, CLCC, MC-Net+, and CDMA methods even fail to locate polyps.
-Other methods only locate some fragments of polyps, while our method accurately and completely segments polyps.
-In the $3^{rd}$ and $4^{th}$ rows, the polyps are visually embedded in their surrounding mucosa, thus it is very difficult to accurately locate and segment these polyps.
+It can be seen that DAN, URPC, CLCC, and SLC-Net methods fail to locate polyps.
+Other methods produce over-segmented fragments and confuse the edges of polyps, while our method accurately and completely segments polyps.
+In the $3^{rd}$ row, the polyps are visually embedded in their surrounding mucosa, thus it is very difficult to accurately locate and segment these polyps.
 From the results, our method performs better than other comparison methods to accurately segment polyps.
-In the $5^{th}$ and $6^{th}$ rows, the polyps have relatively large sizes, making it is challenging to complete locate the polyps.
-In this case, some methods (e.g., URPC, Duo-SegNet, MC-Net+, and BCP) produce several regions with over-segmented fragments, while our method can obtain promising segmentation results and produce fine details of the boundary.
-This is mainly because our method makes use of multi-scale information and cross-generative consistency to learn more powerful feature representations, which are helpful to boost the segmentation performance.
-In addition, we also show some segmentation results containing multiple polyps (see the $7^{th}$ and $8^{th}$ rows).
+In the $4^{th}$ row, the polyps have relatively large sizes, making it challenging to complete locate the polyps.
+In this case, some methods (e.g., UA-MT, URPC, CLCC, CDMA, and MCF) only locate some fragments of polyps, while our method can obtain promising segmentation results and produce fine details of the boundary.
+This is mainly because our method makes use of multi-scale information and cross-generative consistency to learn more powerful feature representations, which help boost the segmentation performance.
+In addition, we also show some segmentation results containing multiple polyps (see the $5^{th}$ row).
 Compared with other methods, the segmentation of multiple polyps by our method is more complete and accurate.
 It can be also observed that our method can effectively locate and segment polyps under different challenging factors, such as scale variation, homogeneous regions, non-sharp boundaries, and multiple polyps.
-TABLE III: Ablative results on the Kvasir and CVC-ColonDB datasets with $30\%$ labeled data.
+[…]
+Figure 7: Qualitative results of our model and other ten compared semi-supervised segmentation methods on brain tumor segmentation dataset.
+[…]
+Figure 8: Qualitative results of our model and other ten compared semi-supervised segmentation methods on skin lesion segmentation dataset.
+IV-C2 Experiments on Brain Tumor Segmentation
+As shown in Tab.
+III, our DEC-Seg achieves significant improvement in five metrics, and compared to the second best, our method achieves $5.2\%$ and $5.5\%$ improvements in terms of mDice and mIou with $10\%$ labeled data, and the corresponding improvements are $2.8\%$ and $4.8\%$ with $30\%$ labeled data.
+Moreover, it can be seen in Fig.
+7 that our method can segment the glioma completely and coherently.
+In the $2^{rd}$ row, the segmentation maps of other methods are discrete, while our method produces more coherent segmented fragments.
+In addition, in the $4^{th}$ row, some methods (e.g., URPC, CLCC, and SLC-Net) do not find glioma regions, while our method accurately locates and segments glioma regions.
+IV-C3 Experiments on Skin Lesion Segmentation
+Tab.
+IV shows the performance on skin lesion segmentation dataset with $10\%$ and $30\%$ labeled ratios.
+DEC-Seg surpasses all state-of-the-arts.
+Our method achieves $1.3\%$ and $3.0\%$ improvements over MCF in terms of mDice and mIou with $10\%$ labeled data, and the corresponding improvements are $1.6\%$ and $3.0\%$ with $30\%$ labeled data.
+Besides, Fig.
+8 shows the visual segmentation results in the skin lesion segmentation dataset.
+Compared with other methods, our DEC-Seg achieves better and more accurate segmentation in blurred boundaries.
+TABLE V: Ablative results on the Kvasir and CVC-ColonDB datasets with $30\%$ labeled data.
 Settings Kvasir CVC-ColonDB
 mDice mIoU $F_{\beta}^{w}$ $S_{\alpha}$ mDice mIoU $F_{\beta}^{w}$ $S_{\alpha}$
-(No.1) Baseline 0.843 0.782 0.829 0.877 0.551 0.489 0.543 0.732
-(No.2) Baseline + SC 0.859 0.801 0.850 0.886 0.600 0.538 0.593 0.757
-(No.3) Baseline + SC + SPC 0.865 0.810 0.856 0.891 0.611 0.548 0.602 0.765
-(No.4) Baseline + SC + SPC + CC 0.875 0.815 0.864 0.892 0.635 0.565 0.627 0.773
-(No.5) Baseline + SC + SPC + CC + CFA 0.876 0.815 0.867 0.892 0.654 0.580 0.649 0.778
-(No.6) Baseline + SC + SPC + CC + CFA + DCF (ours) 0.882 0.822 0.871 0.899 0.691 0.616 0.678 0.798
-IV-C Ablation Study
+(No.1) Baseline 0.850 0.778 0.829 0.873 0.612 0.539 0.600 0.760
+(No.2) Baseline + SC 0.873 0.812 0.865 0.894 0.637 0.572 0.632 0.773
+(No.3) Baseline + SC + DCF 0.882 0.818 0.874 0.892 0.675 0.599 0.658 0.790
+(No.4) Baseline + SC + DCF + CC 0.888 0.827 0.882 0.900 0.702 0.626 0.694 0.807
+(No.5) Baseline + SC + DCF + CFA 0.888 0.824 0.878 0.898 0.698 0.621 0.688 0.804
+(No.6) Baseline + SC + DCF + CC + CFA (ours) 0.893 0.830 0.886 0.902 0.721 0.640 0.709 0.814
+IV-D Ablation Study
 To verify the effectiveness of each key component in DEC-Seg, we conduct ablation studies with $30\%$ labeled data on Kvasir and CVC-ColonDB datasets.
 The ablative results are shown in Tab.
-III, where “Baseline” denotes the base segmentation framework only using the labeled data.
+V, where “Baseline” denotes the semi-supervised framework with only scale-aware perturbation consistency.
 Effectiveness of SC.
-To investigate the importance of scale-enhanced consistency (SC), we add the DC loss to encourage the consistency of predictions from the same inputs with different scales.
+To investigate the importance of scale-enhanced consistency (SC), we add the SC loss to encourage the consistency of predictions from the same inputs with different scales.
 From Tab.
-III, we observe that No.2 (Baseline + SC) outperforms No.1 and obtains an $8.9\%$ improvement in mean Dice.
-This result indicates that scale-enhanced consistency is very helpful to improve polyp segmentation performance and is robust to scale variation.
-Effectiveness of SPC and CC.
-We further study the contributions of scale-aware perturbation consistency (SPC) and cross-generative consistency (CC).
-As shown in Tab.
-III, No.3 performs better than No.2, indicating the effectiveness of scale-aware perturbation consistency.
-Besides, it can be seen that No.4 improves the No.3 performance on the Kvasir dataset, as the mean Dice is improved from $0.865$ to $0.875$.
-Therefore, these improvements indicate that introducing cross-generative consistency loss can help accurately segment polyp tissues in the learning of details and textures.
-Effectiveness of CFA: We then investigate the importance of the proposed CFA module.
-To achieve this, we add the CFA on the basis of No.4, by aggregating the features of adjacent layers in the encoder before feeding into the decoder.
-As reported in Tab.
-III, it can be clearly seen that No.5 has a significant improvement over No.4 on the CVC-ColonDB dataset with a large scale variation, which proves that CFA can effectively capture scale information and improve segmentation performance.
+V, we observe that No.2 (Baseline + SC) outperforms No.1 and obtains a $4.1\%$ improvement in mean Dice.
+This result indicates that scale-enhanced consistency is very helpful in improving polyp segmentation performance and is robust to scale variation.
 Effectiveness of DCF.
 As shown in Tab.
-III, No.6 (using the proposed DCF module) outperforms No.5 on two datasets.
-This indicates that fusion of multi-scale features can further improve the segmentation performance.
+V, No.3 (using the proposed DCF module) outperforms No.2 on two datasets.
+This indicates that the fusion of multi-scale features can further improve the segmentation performance.
 In addition, to further validate the effectiveness of the dual-scale complement fusion strategy in the proposed DCF module, we construct a “Basic” strategy, which conducts a concatenation operation followed by two convolution layers to integrate the two features from different scales.
 The comparison results are shown in Tab.
-IV.
+VI.
 From Tab.
-IV, it can be observed that our DCF performs better than the “Basic” strategy, indicating the effectiveness of the designed dual-scale complementary fusion module.
-Moreover, we visualize the segmentation results by using three different decoders, i.e., the scale-specific and fused decoders, and the comparison results are shown in Fig.
-7.
+VI, it can be observed that our DCF performs better than the “Basic” strategy, indicating the effectiveness of the designed dual-scale complementary fusion module.
+Moreover, we visualize the segmentation results by using three different decoders, i.e., the scale-specific and scale-fused decoders, and the comparison results are shown in Fig.
+9.
 It can be observed that our method can not accurately locate the boundaries of the polyps when only using the original scale features or the downsampled scale features.
-However, we integrate the features from the two scales and then propagate them into the fused decoder, which can produce more accurate segmentation maps (as shown in Fig.
+However, we integrate the features from the two scales and then propagate them into the scale-fused decoder, which can produce more accurate segmentation maps (as shown in Fig.
 9 (c)).
-This further validates that the fusion of the features from different scales is helpful for improving the segmentation performance.
-TABLE IV: Ablative results of our DCF module and a basic fusion strategy on the Kvasir and CVC-ColonDB datasets. The labeled ratio is set to $30\%$.
+This further confirms that integrating features from different scales enhances segmentation performance.
+Effectiveness of CC.
+We further study the contributions of cross-generative consistency (CC).
+As shown in Tab.
+V, it can be seen that No.4 improves the No.3 performance on the CVC-ColonDB dataset, as the mean Dice is improved from $0.675$ to $0.702$.
+Therefore, these improvements indicate that introducing cross-generative consistency loss can help accurately segment polyp tissues in the learning of details and textures.
+Effectiveness of CFA.
+We then examine the significance of the proposed CFA module.
+To do this, we integrate the CFA into configurations No.3 and No.4 by aggregating the features of adjacent layers in the encoder before passing them to the decoder.
+As shown in Tab.
+V, No.5 outperforms No.3, and No.6 shows a substantial improvement over No.4 on the CVC-ColonDB dataset, which has significant scale variation.
+This demonstrates that the CFA module effectively captures scale information and enhances segmentation performance.
+TABLE VI: Ablative results of our DCF module and a basic fusion strategy on the Kvasir and CVC-ColonDB datasets. The labeled ratio is set to $30\%$.
 Settings Kvasir CVC-ColonDB
 mDice mIoU $F_{\beta}^{w}$ mDice mIoU $F_{\beta}^{w}$
-Basic 0.878 0.810 0.866 0.684 0.604 0.676
-DCF 0.882 0.822 0.871 0.691 0.616 0.678
+Basic 0.885 0.823 0.878 0.710 0.633 0.703
+DCF 0.893 0.830 0.886 0.721 0.640 0.709
 […]
 Figure 9: Visualization results of the three decoders.
-(a) predictions of the decoder (i.e., $D_{s1}$, $D_{t1}$) using the original scale features; (b) predictions of the decoder (i.e., $D_{s2}$, $D_{t2}$) using the downsampled scale features; (c) predictions of the fused decoders (i.e., $D_{s}^{f}$, $D_{t}^{f}$) using the multi-scale integrated features.
+(a) predictions of the decoder (i.e., $D_{1}$) using the original scale features; (b) predictions of the decoder (i.e., $D_{2}$) using the downsampled scale features; (c) predictions of the scale-fused decoders (i.e., $D_{f}$) using the multi-scale integrated features.
 The red and blue lines denote the ground truth and predictions, respectively.
+[…]
+Figure 10: Some failure cases generated by our DEC-Seg.
+IV-E Limitation and Future Work
+The qualitative and quantitative evaluations confirm the effectiveness and superiority of our model.
+However, DEC-Seg encounters challenges in accurately segmenting polyps when dealing with ambiguous areas and unclear boundaries.
+Some failure cases generated by our model are depicted in Fig.
+10.
+In the $1^{st}$ and $2^{nd}$ rows, while the prominent polyp regions are segmented accurately, there are inaccuracies in parts areas of the segmentation.
+In the $3^{rd}$ and $4^{th}$ rows, it is evident that the polyp regions are notably small and masked within the background, featuring very unclear boundaries, which presents a significant challenge in accurately segmenting the polyp regions.
+It is apparent that under these circumstances, our DEC-Seg struggles to identify and segment the polyps.
+Therefore, dealing with ambiguous areas and segmenting polyps with unclear boundaries is desired to be investigated in future work.
+Additionally, unlike a simple UNet framework, our model may have higher complexity due to incorporating three decoders and two generative networks.
+Furthermore, our model has only been validated on 2D segmentation tasks.
+Moving forward, we intend to address and overcome these limitations in our future research endeavors.
 V Conclusion
-In this paper, we have presented a novel semi-supervised learning framework (DEC-Seg) for polyp segmentation with colonoscopy images.
+In this paper, we have presented a novel semi-supervised learning framework (DEC-Seg) for medical image segmentation.
 The proposed cross-level feature aggregation module integrates the adjacent features from different resolutions, to enhance the features’ representation ability.
 Then, a scale-enhanced consistency is proposed to handle scale variation and learn more scale-aware features.
-Meanwhile, we design the fused decoders and a dual-scale complementary fusion module to aggregate the features from the scale-specific decoders.
+Meanwhile, we design the scale-fused decoders and a dual-scale complementary fusion module to aggregate the features from the scale-specific decoders and produce the final segmentation maps.
 Moreover, multiple consistency strategies, i.e., scale-aware perturbation consistency and cross-generative consistency, are presented to enhance the learning process and fully leverage unlabeled data to boost the segmentation performance.
-Further, we propose the dual-scale complementary fusion module to integrate the features from two scale-specific decoders, which are passed through the fused decoder to produce the final segmentation maps.
-Experimental results on the five benchmark datasets show that our DEC-Seg is superior to state-of-the-art semi-supervised segmentation methods.
+Experimental results on multiple datasets from three medical image segmentation tasks show that our DEC-Seg is superior to state-of-the-art semi-supervised segmentation methods.